### v2

# Targeted Data Fusion for Causal Survival Analysis Under Distribution Shift

Abstract
Causal inference across multiple data sources offers a promising avenue to enhance the generalizability and replicability of scientific findings. However, data integration methods for time-to-event outcomes, common in biomedical research, are underdeveloped. Existing approaches focus on binary or continuous outcomes but fail to address the unique challenges of survival analysis, such as censoring and the integration of discrete and continuous time. To bridge this gap, we propose two novel methods for estimating target site-specific causal effects in multi-source settings. First, we develop a semiparametric efficient estimator for settings where individual-level data can be shared across sites. Second, we introduce a federated learning framework designed for privacy-constrained environments, which dynamically reweights source-specific contributions to account for discrepancies with the target population. Both methods leverage flexible, nonparametric machine learning models to improve robustness and efficiency. We illustrate the utility of our approaches through simulation studies and an application to multi-site randomized trials of monoclonal neutralizing antibodies for HIV-1 prevention, conducted among cisgender men and transgender persons in the United States, Brazil, Peru, and Switzerland, as well as among women in sub-Saharan Africa. Our findings underscore the potential of these methods to enable efficient, privacy-preserving causal inference for time-to-event outcomes under distribution shift.
1 Introduction
Data fusion, or data integration, promises to improve the generalizability, transportability, and replicability of scientific findings. However, leveraging knowledge from multiple data sources is challenging because of heterogeneous data structures and distribution shift. Furthermore, privacy regulations, such as the General Data Protection Regulation (GDPR) in Europe and the Health Insurance Portability and Accountability Act (HIPAA) in the United States, often preclude pooling individual-level patient data.
Despite rapid growth in data integration methods for causal inference [25, 26, 52, 35, 27, 33], most approaches have been developed for binary or continuous outcomes [11, 5, 16], which do not capture the timing of events. In many clinical and policy-relevant settings, the timing of an outcome—such as time to hospitalization, relapse, or death—is often as important as the occurrence itself. Methods that ignore this temporal dimension risk overlooking treatment effect heterogeneity over time and can limit their utility for decision-making by patients, clinicians, and policymakers. While some recent work has extended data fusion techniques to survival outcomes, these methods impose strong parametric or structural assumptions, such as the Cox proportional hazards (PH) model [29, 23] or a common conditional outcome distribution (CCOD) across sites [32, 7]. Violations of PH or heterogeneous outcome models can result in biased estimates of treatment effects. Moreover, assumptions of strict exchangeability across sources—frequently adopted in the data fusion literature—may fail in the presence of covariate shift, heterogeneous censoring mechanisms, or other types of distribution shift.
Traditionally, time-to-event outcomes have been analyzed using data from single-site randomized clinical trials, with survival curves estimated via the unadjusted Kaplan–Meier estimator [31]. In studies with rich covariate information, extensions have included semiparametric models like the Cox PH model [14, 51, 6] and doubly robust estimators [3]. More recently, Westling et al. [48] integrated double/debiased machine learning (DML) [8] to leverage flexible nuisance estimators [50, 15, 45] to estimate survival curves. However, this approach was designed for single-study settings, and it remains unclear how to generalize the methodology to multi-source contexts, particularly when data sharing is restricted by privacy constraints.
1.1 Our contributions
In this paper, we propose two methods for multi-source causal survival analysis that accommodate events occurring in either continuous or discrete time. First, in settings where individual-level data can be shared and a CCOD assumption is adopted, we derive an estimator based on the semiparametric efficient influence function (EIF) (Section 2.2), which adjusts for covariate shift across sites while achieving both consistency and improved efficiency. Notably, our approach does not require the treatment or censoring mechanisms to be homogeneous across sources; nonetheless, we show that these nuisance functions can be estimated using pooled (global) models without sacrificing asymptotic guarantees. We establish uniform consistency and uniform asymptotic normality of the proposed estimator, thereby providing rigorous theoretical support for its use in multi-source survival settings.
Second, in more restrictive settings where individual-level data sharing is not allowed and the CCOD assumption may not hold, we introduce a federated learning framework (Section 2.3) that leverages a data-adaptive weighting scheme. This approach reweights source sites by up-weighting those more aligned with the target population and down-weighting—or even outright excluding—those likely to introduce bias, all without requiring access to individual-level data. Under a working partial CCOD assumption, we derive site-specific EIF-based estimators and show that all necessary quantities can be computed locally at each site, save for a density ratio term that can be consistently estimated using coarsened summary statistics. To aggregate the local estimators, we formulate and solve a constrained convex optimization problem to obtain federated weights that adaptively emphasizing informative sources while attenuating the influence of less relevant or biased sites. We establish the consistency and asymptotic normality of the proposed federated estimator and provide sufficient conditions under which it achieves efficiency gains over target-only estimation, i.e. its asymptotic variance is strictly smaller than that of an estimator relying solely on the target site.
For both approaches, we develop practical estimators that integrate cross-fitting [8] and off-the-shelf ensemble machine learning methods for nuisance function estimation [45, 17]. By leveraging flexible machine learning tools, we avoid reliance on restrictive parametric or semiparametric assumptions, while still achieving parametric convergence rates under standard regularity conditions.
We assess the empirical performance of our methods through comprehensive simulation studies and an application to the Antibody Mediated Protection (AMP) trials–a set of parallel, multi-site randomized controlled trials conducted across the Americas, Europe, and sub-Saharan Africa to evaluate the efficacy of the broadly neutralizing antibody VRC01 in preventing HIV-1 acquisition [12]. Our results demonstrate that the proposed estimators enable valid, efficient, and privacy-preserving causal effect estimation in complex multi-source survival settings, even under distributional shifts.
2 Multi-Source Inference on Treatment-specific Survival Functions
2.1 Notation, estimand, and assumptions
Consider $K$ studies, each of which could be randomized or observational. For each participant in a given study, we observe a covariate vector $\mathbf{X}$ representing baseline characteristics, and a binary treatment variable $A\in\{0,1\}$, where 0 denotes control and 1 denotes the active treatment. Under the Neyman-Rubin potential outcomes framework [38, 42], we assume each participant has two potential event times, $T^{(1)}$ and $T^{(0)}$, and two potential censoring times, $C^{(1)}$ and $C^{(0)}$, under treatments $A=1$ and $A=0$, respectively. Under the stable-unit treatment value assumption (SUTVA) [41], we have $T=T^{(A)}=AT^{(1)}+(1-A)T^{(0)}$ and $C=C^{(A)}=AC^{(1)}+(1-A)C^{(0)}$ for each participant; under right-censoring, we observe only the possibly censored outcome $Y=\min(T,C)$, together with an event indicator $\Delta=\mathbb{I}(T\leq C)$, where $\mathbb{I}(\cdot)$ is the indicator function. We denote an arbitrary observation as $\mathcal{O}$ and the observed data of sample size $n$ from all study sites as $\{\mathcal{O}_{i}=(\mathbf{X}_{i},A_{i},Y_{i},\Delta_{i},R_{i}),i=1,\dots,n\}$, where $R$ is a site variable taking values in $\{0,1,\dots,K-1\}$, with 0 representing the target site and the remaining $K-1$ values representing source sites, $K>1$.
In this paper, our estimand of interest is the target–site, treatment–specific survival function:
$\theta^{0}(t,a)=\mathbb{P}(T^{(a)}>t\mid R=0),\quad a\in\{0,1\},\quad t\in[0,\tau],$
for some positive $\tau<\infty$. Here, the upper subscript 0 of $\theta^{0}(t,a)$ indicates the target site $R=0$, and $\mathbb{P}$ denotes the distribution of an ideal data unit.
[…]
is the conditional cumulative hazard function under $R=k$, with $N^{k}_{\delta}(t\mid a,\mathbf{x})=\mathbb{P}(Y\leq t,\Delta=\delta\mid A=a,\mathbf{X}=\mathbf{x},R=k)$ the conditional cumulative incidence function of event $(\delta=1)$ or censoring $(\delta=0)$, and $D^{k}(t\mid a,\mathbf{x})=\mathbb{P}(Y\geq t\mid A=a,\mathbf{X}=\mathbf{x},R=k)$. Under discrete time the product integral reduces to an ordinary product $\prod$, and under continuous time to $\exp\{-\Lambda^{k}(t\mid a,\mathbf{x})\}$.
Throughout, we make the following assumptions.
Assumption 2.1 (Unconfoundedness).
$A\perp\mkern-10.0mu\perp T^{(a)}\mid\mathbf{X},R$ and $A\perp\mkern-10.0mu\perp C^{(a)}\mid\mathbf{X},R$.
Assumption 2.2 (Treatment-specific non-informative censoring).
$C^{(a)}\perp\mkern-10.0mu\perp T^{(a)}\mid A=a,\mathbf{X},R$.
Assumption 2.3 (Positivity).
There exists an $\eta\in(0,\infty)$, such that $\mathbb{P}(R=k)\geq 1/\eta$ and for $\mathbb{P}$-almost all $\mathbf{x}$, $\min\cup_{k=0}^{K-1}\{\pi^{k}(a\mid\mathbf{x}),G^{k}(t\mid a,\mathbf{x})\}\geq 1/\eta$ and $\min_{k}S^{k}(t\mid a,\mathbf{x})>0$. Here, $\pi^{k}(a\mid\mathbf{x})=\mathbb{P}(A=a\mid\mathbf{X}=\mathbf{x},R=k)$ is the propensity score of the treatment $A=a$ and $G^{k}(t\mid a,\mathbf{x})=\mathbb{P}(C>t\mid A=a,\mathbf{X}=\mathbf{x},R=k)$ is the conditional survival function of the censoring time in site $k$, for $k=0,1,\dots,K-1$, $a\in\{0,1\}$ and $t\in[0,\tau]$.
The condition $\mathbb{P}(R=k)\geq 1/\eta$ in Assumption 2.3 ensures that the sampling or survey process prevents any site from contributing only an extremely small proportion of participants [32, 7].
Throughout, we define the following notation $\mathcal{H}_{t,a}$ for convenience:
$\mathcal{H}_{t,a}(\mathcal{O};S,G)=\frac{\mathbb{I}(Y\leq t,\delta=1)}{S(Y\mid a,\mathbf{X})G(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda(du\mid a,\mathbf{X})}{S(u\mid a,\mathbf{X})G(u\mid a,\mathbf{X})}.$ (1)
Since $S$ and $\Lambda$ have a one-to-one mapping, we did not involve $\Lambda$ in $\mathcal{H}_{t,a}(\mathcal{O};S,G)$. If we only use data from the target site $(R=0)$, the EIF of $\theta^{0}(t,a)$ can be given in Proposition 2.4.
Proposition 2.4 (Westling et al. [48]).
The nonparametric EIF for $\theta^{0}(t,a)$ given $t\in[0,\tau]$ and $a\in\{0,1\}$ is $\varphi^{*0}_{t,a}(\mathcal{O};S^{0},G^{0},\pi^{0})=\varphi^{*0}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{0}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where
$\varphi^{0}_{t,a}(\mathcal{O};\mathbb{P})=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\bigg{[}1-\frac{\mathbb{I}(A=a)}{\pi^{0}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};S^{0},G^{0})\bigg{]}S^{0}(t\mid a,\mathbf{X}).$
Throughout, let $\mathbb{P}_{n}[f(\mathcal{O})]=n^{-1}\sum_{i=1}^{n}f(\mathcal{O}_{i})$ be the empirical average and $\widehat{\mathbb{P}}$ denote that the nuisance functions are replaced by their estimates. The above EIF motivates a doubly robust estimator $\widehat{\theta}^{0}_{n}(t,a)=\mathbb{P}_{n}[\widehat{\varphi}^{0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})]$. Under regularity conditions, $\widehat{\theta}^{0}_{n}(t,a)$ is regular and asymptotically linear (RAL) and attains the nonparametric efficiency bound uniformly over $t\in[0,\tau]$.
However, that result uses only data from the target site $R=0$. To further improve precision by borrowing strength from the $K-1$ source sites, we develop two strategies in the next two sections.
2.2 Efficient estimation under CCOD and data sharing
When target and source sites are sufficiently similar, it is often plausible to posit a common conditional outcome distribution (CCOD).
Assumption 2.5 (Common conditional outcome distribution).
$T^{(a)}\perp\mkern-10.0mu\perp R\mid\mathbf{X}$ for $a\in\{0,1\}$.
Assumption 2.5 implies $S^{k}(t\mid a,\mathbf{X})=\bar{S}(t\mid a,\mathbf{X})$ for all sites $k$, while still allowing arbitrary shifts in the covariate distribution. Under this CCOD assumption, data from the source sites can be incorporated to yield more efficient estimators. We begin by extending Proposition 2.4 to this setting. Figure 1 presents the directed acyclic graph (DAG) for our full data structure, highlighting the roles of covariates, treatment, site, outcome, and censoring under CCOD.
Figure 1: DAG for data structures under different CCOD assumptions. When CCOD holds (Panel (a)), $R$ and $T$ are conditionally independent given treatment $A$ and covariates $\mathbf{X}$, consistent with Assumption 2.5. When CCOD is potentially violated, as highlighted by the orange dashed arrow in Panel (b), $R$ and $T$ are not necessarily conditionally independent.
Consider the following “global” nuisance functions (each defined without conditioning on $R$): the global propensity score $\bar{\pi}(a\mid\mathbf{x})=\mathbb{P}(A=a\mid\mathbf{X}=\mathbf{x})$, the global conditional survival function for the event time $\bar{S}(t\mid a,\mathbf{x})=\mathbb{P}(T>t\mid A=a,\mathbf{X}=\mathbf{x})$, and the global conditional survival function for the censoring time $\bar{G}(t\mid a,\mathbf{x})=\mathbb{P}(C>t\mid A=a,\mathbf{X}=\mathbf{x})$. Moreover, let the target site propensity be given by $q^{0}(\mathbf{X})=\mathbb{P}(R=0\mid\mathbf{X})$. Similar to Assumption 2.3, we impose the following positivity assumptions for these nuisance functions.
Assumption 2.6 (Positivity of global nuisance functions).
There exists $\eta\in(0,\infty)$ such that for $\mathbb{P}$-almost all $\mathbf{x}$, $\min\{\bar{\pi}(a\mid\mathbf{x}),\bar{G}(t\mid a,\mathbf{x})\}\geq 1/\eta$ and $\bar{S}(t\mid a,\mathbf{x})>0$, for $a\in\{0,1\}$ and $t\in[0,\tau]$.
Theorem 2.7.
Under Assumption 2.5, the semiparametric EIF of $\theta^{0}(t,a)$ given $t\in[0,\tau]$ and $a\in\{0,1\}$ is $\varphi^{*\text{CCOD}}_{t,a}(\mathcal{O};\bar{S},\bar{G},\bar{\pi})=\varphi^{*\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where
$\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})=\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}-\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};\bar{S},\bar{G})\bigg{]}\bar{S}(t\mid a,\mathbf{X}).$
Given this EIF, we propose a semiparametrically efficient “CCOD estimator:”
$\widehat{\theta}^{\text{CCOD}}_{n}(t,a)=\mathbb{P}_{n}[\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})].$
We employ $M$-fold cross-fitting following the DML framework [8], using flexible ensemble learning via the survSuperLearner [48] and SuperLearner [45] packages to estimate the nuisance functions. In Theorem 2.8, we demonstrate the (uniform) RAL property of the proposed CCOD estimator.
Theorem 2.8.
If Conditions D.1D.3 in Appendix D.1 (i.e., consistency, positivity, and rate condition on the product of prediction errors for the nuisance functions) hold, then $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{t,a})+o_{p}(n^{-1/2})$. In particular, $n^{1/2}(\widehat{\theta}^{\text{CCOD}}_{n}(t,a)-\theta^{0}(t,a))$ convergences in distribution to a normal random variable with mean zero and variance $\sigma^{2}=\mathbb{P}[(\varphi^{*\text{CCOD}}_{t,a})^{2}]$. If Conditions D.4D.5 (which strengthen the consistency and product error conditions to hold uniformly) also hold, then
$\sup_{u\in[0,t]}\left|\widehat{\theta}^{\text{CCOD}}_{n}(u,a)-\theta^{0}(u,a)-\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{u,a})\right|=o_{p}(n^{-1/2}).$
In particular, $\{n^{1/2}(\widehat{\theta}^{\text{CCOD}}_{n}(u,a)-\theta^{0}(u,a)):u\in[0,t]\}$ converges weakly as a process in the space $\ell^{\infty}([0,t])$ of uniformly bounded functions on $[0,t]$ to a tight mean zero Gaussian process with covariance function $(u,v)\mapsto\mathbb{P}(\varphi^{*\text{CCOD}}_{u,a}\varphi^{*\text{CCOD}}_{v,a})$.
Remark 2.9.
Among the global nuisance functions $\bar{S}$, $\bar{\pi}$, and $\bar{G}$, only $\bar{S}$ is assumed (based on CCOD) to equal each corresponding site-specific function (in this case, $S^{k}$, for all $k$). The inclusion of $\bar{\pi}$ and $\bar{G}$ in Theorem 2.7 arises purely from the EIF derivation and does not require homogeneity assumptions on the distributions of $A$ or $C^{(a)}$ across sites. In practice, $A$ and $C^{(a)}$ are modeled as functions of $\mathbf{X}$ using the pooled sample. Moreover, inspecting the forms of the EIFs in Proposition 2.4 and Theorem 2.7, it is clear that the CCOD estimator, unlike the target-only estimator, incorporates data from all sites: while the “anchor” term of $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$ relies on target data only, the “augmentation” term incorporates information from all sites.
2.3 Federated estimation under heterogeneous outcome distributions across sites
In practice, it will often be unreasonable to assume CCOD, and sharing data might not be permitted under certain privacy constraints. When the CCOD assumption is violated, some source sites may provide relevant information for constructing the target-site survival estimate, while some may not. In this section, we present a methodology that combines information from target and source sites in a data-adaptive, privacy-preserving manner. In particular, the approach involves only minimal data sharing of summary statistics across sites.
2.3.1 Local Sites
We first construct robust site-specific survival curve estimates using data from site $k\in\{0,1,\dots,K-1\}$, making a working partial CCOD assumption that outcomes have the same conditional distribution in site $k$ as in the target site when $k\not=0$. Note that we use this working partial CCOD assumption only to derive the form of the EIF, serving as the first step of our algorithm; to aggregate information from source sites, we derive federated weights to account for possible violations of CCOD in Section 2.3.2. An EIF under this assumption is derived in the following theorem.
Theorem 2.10.
For $k\in\{0,1,\dots,K-1\}$, $\theta^{0}(t,a)$ is a pathwise differentiable parameter given $t\in[0,\tau]$ and $a\in\{0,1\}$, with semiparametric EIF—under the working assumption that $S^{k}\equiv S^{0}$— given by $\varphi^{*k,0}_{t,a}(\mathcal{O};S^{k},S^{0},G^{k},\pi^{k},\omega^{k,0})=\varphi^{*k,0}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where
$\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}S^{0}(t\mid a,\mathbf{X})-\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k)}\omega^{k,0}(\mathbf{X})S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};S^{k},G^{k}),$
where $\omega^{k,0}(\mathbf{X})=\dfrac{\mathbb{P}(\mathbf{X}\mid R=0)}{\mathbb{P}(\mathbf{X}\mid R=k)}$ is a density ratio function of covariates $\mathbf{X}$ under target site to source site $k$.
A site-specific estimator can be motivated from the EIF above, namely $\widehat{\theta}^{k,0}_{n}(t,a)=\mathbb{P}_{n}[\widehat{\varphi}^{k,0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})]$. Of note regarding $\widehat{\theta}^{k,0}_{n}(t,a)$: (i) the “anchor” term in the EIF uses target-site data ($\mathbb{I}(R=0)$), while the “augmentation” term leverages site-$k$ data ($\mathbb{I}(R=k)$); (ii) all information is obtained independently from each site except for the density ratio $\omega^{k,0}(\mathbf{X})$, which is estimated by sharing only coarse summary statistics under flexible models [25], thereby adhering to data sharing constraints; and (iii) for $S^{k}$ in the augmentation term, we train an $S^{0}$ model on the target site and apply its predictions to site $k$, since $S^{0}$ and $S^{k}$ are exchangeable under partial CCOD. Crucially, if partial CCOD is violated, we can detect site heterogeneity by the difference between $\widehat{\theta}^{k,0}_{n}(t,a)$ and $\widehat{\theta}^{0}_{n}(t,a)$.
Furthermore, we use $\varphi^{*k}_{t,a}$ and $\varphi^{k}_{t,a}$ to denote the centered and uncentered EIFs for targeting $\theta^{k}(t,a)$ (the survival of site $k$). These follow the form in Proposition 2.4, with the site $R=0$ replaced by $R=k$. In the case where $k=0$, $\varphi^{*0,0}_{t,a}=\varphi^{*0}_{t,a}$ and $\varphi^{0,0}_{t,a}=\varphi^{0}_{t,a}$. The estimator $\widehat{\theta}^{k}(t,a)=\mathbb{P}_{n}[\varphi^{k}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})]$ serves as an “unadjusted” estimator for $\theta^{0}(t,a)$, lacking robustness when $k\not=0$. In comparison, with the anchor term adjusted using target site data, $\widehat{\theta}^{k,0}_{n}(t,a)$ preserves robustness for targeting $\theta^{0}(t,a)$.
Remark 2.11.
To ensure the monotonicity of the estimated site-specific survival curves, we invoke isotonic regression techniques [49], which enforce a non-increasing constraint on the survival estimates, thereby maintaining their logical consistency over time.
2.3.2 Aggregation across sites
To aggregate information from the target and source sites, for a fixed time $t$ and treatment $a$, we use an $L_{1}$-penalization procedure to calculate federated weights for each site. We define $\widehat{\chi}_{n,t,a}^{k,0}=\widehat{\theta}^{k,0}(t,a)-\widehat{\theta}^{0}(t,a)$, and write $\boldsymbol{\eta}_{t,a}=(\eta^{0}_{t,a},\eta^{1}_{t,a},\dots,\eta^{K-1}_{t,a})$ for the vector of federated weights. We estimate $\boldsymbol{\eta}_{t,a}$ data-adaptively by minimizing the following $L_{1}$-penalized objective function:
$Q(\boldsymbol{\eta}_{t,a})=\mathbb{P}_{n}\left[\left\{\widehat{\varphi}^{*0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})-\sum_{k=1}^{K-1}\eta^{k}_{t,a}\widehat{\varphi}^{*k,0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})\right\}^{2}\right]+\frac{1}{n}\lambda\sum_{k=1}^{K-1}|\eta^{k}_{t,a}|(\widehat{\chi}_{n,t,a}^{k,0})^{2},$ (2)
subject to $0\leq\eta^{k}_{t,a}\leq 1$, where $\widehat{\chi}_{n,t,a}^{k,0}=\widehat{\theta}^{k,0}(t,a)-\widehat{\theta}^{0}(t,a)$, for all $k=0,1,\dots,K-1$ and $\sum_{k=0}^{K-1}\eta^{k}_{t,a}=1$, and $\lambda$ is a tuning parameter that controls the bias-variance trade-off and is chosen by cross-validation.
The objective function follows the principle of assigning higher (resp. lower) weight to more (resp. less) informative sites. Heuristically, our approach anchors on the nonparametric estimate $\widehat{\theta}^{0}_{n}(t,a)$ and weights site $k$ when it is deemed similar to the target site. We then compute $\widehat{\theta}^{\text{fed}}_{n}(t,a)$, our “federated” (FED) estimator, as the weighted average of the site-specific survival estimates.
$\widehat{\theta}^{\text{fed}}_{n}(t,a)=\sum_{k=0}^{K-1}\widehat{\eta}^{k}_{t,a}\widehat{\theta}^{k,0}_{n}(t,a).$
Figure 2 below gives an intuitive visualization of the proposed FED algorithm. We also present all details for the FED method in Algorithm 1 in Appendix C.
Figure 2: Illustration of the Federated Algorithm. Each site has its underlying survival functions, which are first estimated in a site-specific manner. Then, for all $a\in\{0,1\}$ and $t\in[0,\tau]$, treatment- and time-specific federated weights for the target and source sites are derived, ultimately producing the weighted average survival functions (the purple step).
2.3.3 Theoretical results of the federated estimator
We establish the asymptotic properties and efficiency gain of the federated estimator in the following result.
Theorem 2.12.
If Conditions D.6D.8 in Appendix D.2 (i.e., consistency, positivity, and certain rate conditions on the product of prediction errors of the nuisance functions at each local site) and oracle selection conditions for the federated weights $\boldsymbol{\eta}_{t,a}$ in Appendix D.3 hold, then the federated estimator $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ at each $(t,a)\in[0,\tau]\times\{0,1\}$ has asymptotic distribution
$\sqrt{n/\widehat{\mathcal{V}}_{t,a}}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a)-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,1),$
where $\widehat{\mathcal{V}}_{t,a}$ is a consistent estimator for the asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ (see Appendix D.3 for its form), which is no greater than the asymptotic variance of the target-only estimator, $\widehat{\theta}^{0}_{n}(t,a)$. Moreover, if any source site provides a consistent estimate of $\theta^{0}(t,a)$, then the asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ is strictly smaller than that of $\widehat{\theta}^{0}_{n}(t,a)$.
All technical proofs are presented in Appendices D.2 and D.3. Here, we briefly outline the key theoretical insights and their justifications. To quantify the efficiency gain afforded by our federated method, we adopt an approach analogous to that in existing federated learning studies [25, 33, 27].
Letting $\mathcal{S}=\{1,\dots,K-1\}$ denote the set of source sites, we define the oracle selection space for $\boldsymbol{\eta}_{t,a}$ and the corresponding weight space as, respectively, $\mathcal{S}^{*}_{t,a}=\{k\in\mathcal{S}:\theta^{k}(t,a)=\theta^{0}(t,a)\}$ and $\mathbb{R}^{S^{*}_{t,a}}=\{\boldsymbol{\eta}_{t,a}\in\mathbb{R}^{K-1}:\eta^{j}_{t,a}=0,\forall j\not\in\mathcal{S}^{*}_{t,a}\}$. Appendix D.3 delineates conditions under which our federated approach can recover the following optimal weights by solving (2):
$\bar{\boldsymbol{\eta}}_{t,a}=\underset{\eta^{k}_{t,a}=0,\forall k\not\in\mathcal{S}^{*}_{t,a}}{\text{arg min}}\mathcal{V}_{t,a}(\boldsymbol{\eta}_{t,a}),$
where $\mathcal{V}_{t,a}(\boldsymbol{\eta}_{t,a})$ is the asymptotic variance of the federated estimator under a fixed weight vector $\boldsymbol{\eta}_{t,a}$. Since the target-only estimator corresponds to a special case $\boldsymbol{\eta}_{t,a}=(1,0,\dots,0)$ (assigning weight 1 only to the target site), its asymptotic variance is no larger than that of any federated estimator. Moreover, the key rationale behind the federated weighting is that if the bias term $\widehat{\chi}_{n,t,a}^{k,0}$ remains asymptotically non-zero, the corresponding weight $\eta^{k}_{t,a}$ converges to zero.
3 Numerical Experiments
We conduct simulations to evaluate our “CCOD” and “FED” estimators against competing methods: target-site-only (TGT), simple pooling (POOL), and inverse variance weighting (IVW). TGT uses only target site data ($R=0$), POOL pools data from all sites without adjustment, and IVW computes a weighted average of site-specific estimators, with weights based on inverse variance. We also consider “FED (BOOT),” as another version of FED, which applies bootstrap (200 replicates) when deriving the federated weights using (2), aiming to reduce uncertainty associated with the estimated influence functions.
3.1 Data generation and performance metrics
We conduct $M=500$ Monte Carlo replications, with $n=\sum_{k=1}^{K}n_{k}$ observations distributed across $K=5$ sites. Each site $k\in\{0,1,\dots,4\}$ has $n_{k}$ observations, with site $k=0$ as the target site. We fix the target site sample size at $n_{0}=300$ and vary $n_{k}\in\{300,600,1000\}$ for all $k=1,\dots,4$, corresponding to limited, moderate, and abundant source data. Covariate, treatment, and outcome generation details are provided in Appendix A.1. The true survival curves are derived by averaging over a simulated population of size $n_{\text{super}}=10^{8}$ generated from the target site distribution.
We model time-to-event outcomes over one year (365 days), truncating censoring at day 200. Results are evaluated at days 30, 60, and 90. We introduce site-specific heterogeneity across five cases: (i) homogeneous; (ii) covariate shift; (iii) outcome shift; (iv) censoring shift; and (v) all shifts. Case (i) assumes identical processes across sites, while cases (ii)–(iv) introduce site-specific shifts in covariates, conditional outcome, or conditional censoring mechanism. Case (v) combines all three shifts. Figure 5 in Appendix A.1 illustrates the true underlying survival curves under covariate and outcome shifts, providing a direct comparison of these two types of site heterogeneity.
Simulation results are evaluated using: (A) Estimation Bias, displayed by boxplots from 500 simulations. Narrower boxes and fewer outliers indicate higher efficiency. (B) Relative Root Mean Square Error (RRMSE), calculated as the ratio of a method’s RMSE to the TGT estimator’s RMSE. RRMSE values below 1 indicates efficiency gains, while values above 1 indicate losses. (C) 95% Confidence Interval (CI) Width, which represents the size of the estimated variance. (D) Coverage Probability (CP%), defined by the proportion of replications where the true parameter falls within the CI. Values close to the nominal 95% level reflect accurate variance estimation.
3.2 Results
For succinctness, we present the simulation results only under day 90 and $n_{k}=600$ ($k\geq 1$) in this section. The full simulation results can be found in Appendix A.2. The main results (Figure 3) show that FED and FED (BOOT) consistently exhibit biases centered around zero (panel (A)). Across all site heterogeneity and treatment scenarios, TGT proves the least efficient, while FED and FED (BOOT) achieve a 5%–16% lower RMSE by the RRMSE in panel (B), confirming that our federated methods are consistent with improved efficiency. Furthermore, both federated methods provide 95% CIs that are no wider than TGT’s and produce CPvalues near the nominal 95% level. In addition, Figure 6 in Appendix A.2 displays the federated weights across different scenarios. The figure includes a panel showing average weights over time across 500 simulations, as well as a panel with scatterplots of weights versus $(\widehat{\chi}_{n,t,a}^{k,0})^{2}$. These plots reveal that federated weights generally decrease as $(\widehat{\chi}_{n,t,a}^{k,0})^{2}$ increases. The target site often receives higher weights when covariate and outcome shifts are present. These weights can also vary at different time points based on the similarity of survival function values at a given time.
Figure 3: Simulation results evaluated at day 90, under $n_{0}=300$ and $n_{k}=600$, $k=1,2,3,4$.
In contrast, although POOL and IVW yield smaller variance (as shown by narrower boxplots and CIs), they are highly biased under Covariate, Outcome, and All Shifts, resulting in elevated RRMSE and severely low CP%. However, these methods remain robust under Censoring Shift, consistent with the notion that censoring is estimated separately on each site and plays an ancillary role.
Lastly, the CCOD estimator stands out for its consistency and efficiency under Covariate Shift, aligning with its theoretical strengths (consistent, semiparametrically efficient, and accommodating covariate shift), but it can fail under Outcome or All Shifts due to it reliance on the CCOD assumption.
4 Real Data Analysis
We applied our methods to two coordinated randomized trials, HVTN 704/HPTN 085 and HVTN 703/HPTN 081 [12, 39], which together enrolled 4,611 participants to evaluate whether a broadly neutralizing monoclonal antibody (bnAb) reduces the risk of HIV-1 acquisition. The endpoint was an HIV diagnosis by week 80–a rare event with a 3.77% incidence. The overall rate of loss to follow-up in this study is low (less than 10% in each treatment group), as indicated by Corey et al. [12].
We divided the data into four subsets by region: (i) SA: South Africa, (ii) OA: other sub-Saharan African countries, (iii) BP: Brazil or Peru, and (iv) US: the United States or Switzerland. Participants in (i) and (ii) were women, whereas those in (iii) and (iv) were cisgender men and transgender individuals, potentially introducing important distribution shifts across regions. Because HIV diagnoses were infrequent, we only used 2-fold cross-fitting, and estimated conditional survival for both event and censoring processes by an ensemble of Kaplan–Meier, Cox regression and survival random forest models via the survSuperLearner package [48]. The propensity score and density ratio (used in federated method) models were fitted by the ensemble of logistic regression and LASSO using the Super Learner [45]. In all these nuisance models, age (year), a standardized risk score built by machine learning, and weight (kg) at baseline were included as covariates.
Figure 4: Data analysis results (South Africa as the target site). Panel (A): Missing values early in IVW indicate large IVW weights. Relative efficiency, defined as the ratio of the estimated standard deviation to that of the TGT estimator, presented for three time points: 148, 330, and 512 days. Panel (B): Time-specific federated weights and smoothed weights using locally weighted regression [10].
In this section, we focus on the SA region as our target site (Figure 4). Additional results—including analyses treating other regions as the target (Figure 11), direct comparisons of site-specific survival curves (Figure 10), and summaries of baseline covariates and event rates by region and treatment group (Table 1)—are provided in Appendix B. Table 1 shows that OA’s summary statistics closely mirror those of SA, whereas BP and US differ markedly in mean standardized risk scores, weight, and HIV prevalence, indicating covariate—and possibly outcome—shifts relative to SA. This pattern aligns with panel (B), where SA overall receives the highest federated weights over time, followed by OA, US, and BP.
Panel (A) illustrates that the TGT, FED, and FED (BOOT) methods yield nearly identical survival curves. However, TGT occasionally has wider confidence intervals—reflecting lower efficiency—and fails to produce interval estimates at some early time points when its variance estimates are unavailable or too small. In contrast, FED and FED (BOOT) still provide interval estimates in those cases. The CCOD method generates comparable curves with narrower confidence intervals and lower variance. These efficiency gains for FED, FED (BOOT), and CCOD align with our simulation results in Section 3. By comparison, the IVW and POOL methods—while exhibiting higher nominal efficiency (lower relative-efficiency values)—diverge from the other methods’ survival trends, suggesting potential bias arising from covariate and outcome distribution shifts when targeting the SA population.
5 Discussion
We proposed two methods for estimating treatment-specific survival functions in a target population, leveraging external source data with possibly shifted covariate and outcome distributions. When multi-source data can be pooled and CCOD holds, our CCOD estimator is consistent, semiparametrically efficient, and handles covariate shifts. If CCOD is violated, the federated estimator remains consistent, offers moderate efficiency gains, and preserves patient privacy.
Limitations of our work present opportunities for future work. First, although our simulations show efficiency gains for the federated method, exploring efficient covariate-adaptive weighting [34] may be a promising future avenue. Second, when data sharing is allowed but CCOD may fail, it is unclear if any method outperforms both target-site-only and federated estimators, as our semiparametric efficiency is based on the CCOD assumption. Third, our time-specific federated weights, while flexible, may produce nonsmooth weights over time, and can be computationally intensive in continuous time. Therefore, future work should aim to more efficiently capture temporal trends and appropriately smooth these weights. Finally, future work should aim to incorporate time-varying covariates, which can improve efficiency by leveraging additional predictive information [40]. To the best of our knowledge, existing methods are limited to settings with discrete survival times [1, 19, 13], and unifying the related methodology for continuous time remains an open problem.
Our framework can potentially be extended to various settings, including surrogate markers [20, 28], conformal inference [35], dynamic treatment regimes [55, 9], and data-driven individual selection from external sources [56, 21]. It may also apply to direct treatment effect modeling for survival outcomes, such as marginal structural models [54] and structural nested models [53], as well as other survival estimands like restricted mean survival times [23, 30], competing risk [36] and truncation [37, 24, 47].
Appendix A Additional Details and Results for Numerical Experiments
A.1 Data generating process
Three covariates $X_{1}$, $X_{2}$, and $X_{3}$ are sampled as transformations of Beta random variables with site-specific parameters:
$X_{1}$ $\sim 33\cdot\text{Beta}(1.1-0.05\gamma(k),1.1+0.2\gamma(k))+9+2\gamma(k),$
$X_{2}$ $\sim 52\cdot\text{Beta}(1.5+(X_{1}+0.5\gamma(k))/20,4+2\gamma(k))+7+2\gamma(k),$
$X_{3}$ $\sim(4+2\gamma(k))\cdot\text{Beta}(1.5+|X_{1}-50+3\gamma(k)|/20,3+0.1\gamma(k)),$
where $\gamma(k)$ represents some function of site $k$, specified later. We then generate the treatment assignment probabilities $\pi(\mathbf{X})$ using the logistic function:
$\text{logit}(\pi(\mathbf{X}))=-1.05+\log\left(1.3+\exp(-12+X_{1}/10)+\exp(-2+X_{2}/12)+\exp(-2+X_{3}/3)\right),$
and treatments $A$ are sampled as $A\sim\text{Bernoulli}(\pi(\mathbf{X})).$
Next, we consider the mechanisms of event and censoring times. The hazard rates for event times and censoring times are given by the following $\exp(h_{t})$ and $\exp(h_{c})$, respectively, where $h_{t}=-5.02+0.1(X_{1}-25)-0.1(X_{2}-25)+0.05(X_{3}-2)+D_{T}(k)\cdot 0.1(X_{2}-25)+A\cdot\delta_{T}(k)\cdot 0.1(X_{1}+X_{2}+X_{3}-50),$ and $h_{c}=-4.87+0.01(X_{1}-25)-0.02(X_{2}-25)+0.01(X_{3}-2)-D_{C}(k)\cdot 0.1(X_{2}-25)+A\cdot\delta_{C}(k)\cdot 0.1(X_{1}+X_{2}+X_{3}-50)$.
Here, $D_{T}(k)$, $D_{C}(k)$, $\delta_{T}(k)$ and $\delta_{C}(k)$ are some site-specific indicators, specified later, for varying the treatment effects and trends of survival curves for different sites. Then, event times and censoring times are sampled as:
$T=\left(-\frac{\log(U_{1})}{\exp(h_{t})\cdot\lambda}\right)^{1/\rho},\quad C=\left(-\frac{\log(U_{2})}{\exp(h_{c})\cdot\lambda}\right)^{1/\rho},$
with $\rho=1.2$, $\lambda=0.6$, and $U_{1},U_{2}\sim\text{Uniform}(0,1)$. This technique follows [2]. Thus, the observed times and event indicators are $Y=\min(T,C),\Delta=\mathbb{I}(T\leq C)$, respectively. Under this data generating process (DGP), the event time is generated as a similar way of days in a year (365 days), and we truncate the censoring time at $t=200$ days to mimic the end of follow-up in survival analysis. Our DGP supports the following scenarios based on site-specific heterogeneity:
  • Homogeneous: Homogeneous covariates and hazard rates across sites. We let $\gamma(k)=D_{T}(k)=D_{C}(k)=\delta_{T}(k)=\delta_{C}(k)=0$ for $k=0,1,\dots,4$.
  • Covariate Shift: Covariates $X_{1}$, $X_{2}$, and $X_{3}$ vary across sites. We let $\gamma(k)=k$ and $D_{T}(k)=D_{C}(k)=\delta_{T}(k)=\delta_{C}(k)=0$, for $k=0,1,\dots,4$.
  • Outcome Shift: Conditional outcome distribution varies across sites. We assign $\gamma(k)=0$, $D_{T}(k)=\delta_{T}(k)=k$, and $D_{C}(k)=\delta_{C}(k)=0$ for $k=0,1,\dots,4$.
  • Censoring Shift: Censoring mechanism varies across sites. We let $\gamma(k)=0$, $D_{T}(k)=\delta_{T}(k)=0$ and $D_{C}(k)=\delta_{C}(k)=k$, for $k=0,1,\dots,4$.
  • All Shift: Covariates and both event and censoring effects vary across sites. We let $\gamma(k)=D_{T}(k)=D_{C}(k)=\delta_{T}(k)=\delta_{C}(k)=k$, for $k=0,1,\dots,4$.
Figure 5 plots the true treatment-specific survival curves under the Covariate Shift and Outcome Shift scenarios, as defined by our designed DGPs, to illustrate the effect of site differences on survival outcomes.
Source sites have covariate shift only: Refer to caption
Source sites have outcome (label) shift only: Refer to caption
Figure 5: True treatment-specific survival curves across different sites. Each curve is derived from a random sample of $n=10^{4}$ generated under the site’s own DGP. Dashed lines represent the target site’s survival curves for reference. Under covariate shift, the curves maintain similar shapes and trends, differing primarily in scale. In contrast, outcome shift leads to marked alterations in the shapes of the survival curves.
A.2 Complete simulation results
Figure 6: Upper panel: Average federated weights of each site at different time point by site heterogeneity cases. Lower panel: catter plots of site-specific federated weights vs. discrepancy measure $(\widehat{\chi}^{k}_{n,t,a})^{2}$ values, under 5 scenarios of site heterogeneity and 3 selected time points (days 30, 60 and 90). Sites 2–4 under Covariate Shift and All Shift have more larger $(\widehat{\chi}^{k}_{n,t,a})^{2}$ values with clear trends of decreasing weights. The pink dashed lines indicate weight $=$ 1/5, i.e., one over five sites. This figure uses the case where $n_{k}=300$ ($k\geq 1$) as an illustration for weights.
Figure 7: Estimation bias (boxplots), relative root mean square error (RRMSE) compared to TGT, coverage probability (CP%) with 95% nominal coverage level, and width of 95% confidence interval (CI) under $n_{k}=300$ ($k\geq 1$), evaluated at days 30, 60 and 90 in simulation.
Figure 8: Estimation bias (boxplots), relative root mean square error (RRMSE) compared to TGT, coverage probability (CP%) with 95% nominal coverage level, and width of 95% confidence interval (CI) under $n_{k}=600$ ($k\geq 1$), evaluated at days 30, 60 and 90 in simulation.
Figure 9: Estimation bias (boxplots), relative root mean square error (RRMSE) compared to TGT, coverage probability (CP%) with 95% nominal coverage level, and width of 95% confidence interval (CI) under $n_{k}=1000$ ($k\geq 1$), evaluated at days 30, 60 and 90 in simulation.
Appendix B Additional Data Analysis Results
Table 1 presents summary statistics for baseline covariates and outcomes in the AMP trial data, stratified by region and treatment group. Comparing the treatment groups—both overall and within each region—we observe that the treated group consistently shows a lower average event proportion. Additionally, some covariates appear to shift across regions; for example, among treated participants, the standardized risk scores exhibit notably different means when comparing SA to BP and US.
Treated (bnAb) group
Total ($n=3,076$) SA ($n=679$) OA ($n=608$) BP ($n=846$) US ($n=943$)
Age (year) at baseline 25.9 (4.60) 27.0 (5.19) 25.4 (4.59) 25.1 (3.70) 26.2 (4.68)
Standardized risk score 0.0 (1.00) -0.01 (1.00) 0.02 (1.00) 0.76 (0.67) -0.68 (0.71)
Weight at baseline (kg) 72.8 (15.64) 68.8 (14.24) 65.2 (12.63) 70.9 (12.42) 82.3 (16.43)
HIV diagnosis by week-80 107 (3.48%) 27 (3.98%) 20 (3.29%) 46 (5.44%) 14 (1.49%)
Control (placebo) group
Total ($n=1,535$) SA ($n=340$) OA ($n=297$) BP ($n=428$) US ($n=470$)
Age (year) at baseline 25.9 (4.72) 26.6 (5.28) 25.4 (4.78) 25.2 (3.94) 26.1 (3.79)
Standardized risk score 0.0 (1.00) 0.02 (0.92) -0.02 (0.98) 0.75 (0.67) -0.68 (0.73)
Weight at baseline (kg) 72.5 (16.35) 67.6 (14.77) 65.1 (13.64) 71.1 (12.84) 81.8 (17.5)
HIV diagnosis by week-80 67 (4.36%) 16 (4.71%) 13 (4.38%) 29 (6.78%) 9 (1.91%)
Table 1: Summary statistics of AMP trial data by treatment group and region. The standardized risk score is a baseline score built by machine learning models [12] that is predictive to the time-to-event outcome. Age, standardized risk score and weight are summarized by mean (standard deviation), while the HIV diagnosis by week-80 is summarized by count (proportion).
In Figure 10, we plot the region-specific survival curves of all the 4 regions we considered (SA, OA, BP and US) for a direct comparison on region heterogeneity, using their target-site-only (TGT) estimators, to showcase the heterogeneous effects of the bnAb antibody treatment on different target populations.
Figure 10: Estimated region-specific survival curves of the HVTN 704/HPTN 085 and HVTN 703/HPTN 081 trials. SA (our target region in the main text) and OA exhibit relatively similar curves, indicating less heterogeneity of these two regions. In contrast, both BP and US regions show significant differences to SA, which also confirms why they often have small or zero federated weights in Panel (B) of Figure 4 in the main text. The BP and US also show a substantial difference on their curves.
Figure 11: Additional data analysis results when treating the other three regions (OA, BP and US) as the target site.
Furthermore, in Figure 11, we present the results—including survival curve estimations and federated weights—using three regions other than SA as the target population. For the federated weights, similar to Figure 4 in the main text, we applied locally weighted regression [10] to smooth the observed weights over the study period, providing a clearer visualization of temporal trends in this specific example.
From Figure 11, we observe that for each region, the FED and FED (BOOT) methods yield results similar to the TGT estimator, while also recovering some interval estimations at earlier time points. This finding is consistent with the observations made in Figure 4. In contrast, the IVW and POOL methods deviate noticeably from the TGT, FED, and FED (BOOT) results—especially for the BP and US regions—indicating potential biases introduced by site heterogeneity. The CCOD in each region produces curves that more closely align with the corresponding TGT, FED, and FED (BOOT) methods.
Finally, regarding federated weights, the results for the OA region resemble those of SA in Figure 4. However, for the BP and US regions, the federated weights are nearly 1 for the target site and 0 for all other sites. This pattern suggests that when targeting the survival curves of BP or US, other sites contribute substantial biases—an observation that corroborates our findings in Figure 10.
Appendix C Details of the Proposed Algorithms
In this section, we describe (i) the procedure for federated estimation in Algorithm 1, and (ii) the DML procedure for fitting and predicting nuisance functions in Algorithm 2. Algorithm 1 implements the method introduced in Section 2.2 that is also illustrated in Figure 2. Algorithm 2 outlines the DML procedure used within Algorithm 1, and a similar approach is applied to the CCOD and other estimators when predicting their respective nuisance functions.
Algorithm 1 Federated Learning for Multi-Source Causal Survival Analysis.
1:  Input: Multi-source right-censored data $\{\mathcal{O}_{i}=(\mathbf{X}_{i},A_{i},Y_{i},\Delta_{i},R_{i}),i=1,\dots,n\}$, a time horizon $\tau>0$, and a fine time grid $\{0,\epsilon,2\epsilon,\dots,\tau\}$ for $[0,\tau]$, $\epsilon>0$.
2:  Output: Estimated treatment-specific survival curves $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ and variance $\widehat{\mathcal{V}}_{t,a}=\widehat{\mathcal{V}}(\widehat{\theta}^{\text{fed}}_{n}(t,a))$ for $a\in\{0,1\}$ and $t\in\{0,\epsilon,2\epsilon,\dots,\tau\}$.
3:  for $(t,a)\in\{0,\epsilon,2\epsilon,\dots,\tau\}\times\{0,1\}$ do
4:     Estimate the EIFs using the DML algorithm with sample-splitting and cross-fitting described in Algorithm 2.
5:     Obtain the target and source estimators as $\widehat{\theta}^{k,0}_{n}(t,a)=\mathbb{P}_{n}\left[\widehat{\varphi}^{k,0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})\right]$, for all $k=0,1,\dots,K-1$.
6:     Obtain the site-specific discrepancy measure (difference of the target and source estimators) as $\widehat{\chi}_{n,t,a}^{k,0}=\mathbb{P}_{n}\left[\widehat{\varphi}^{k,0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})-\widehat{\varphi}^{0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})\right]$, $k=1,\dotsm K-1$.
7:     Solve for aggregation treatment- and time-specific weights $\widehat{\boldsymbol{\eta}}_{t,a}=(\widehat{\eta}^{0}_{t,a},\widehat{\eta}^{1}_{t,a},\dots,\widehat{\eta}^{K-1}_{t,a})$ that minimize
$Q(\boldsymbol{\eta}_{t,a})=\mathbb{P}_{n}\left[\left\{\widehat{\varphi}^{*0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})-\sum_{k=1}^{K-1}\eta^{k}_{t,a}\widehat{\varphi}^{*k,0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})\right\}^{2}\right]+\frac{1}{n}\lambda\sum_{k=1}^{K-1}|\eta^{k}_{t,a}|(\widehat{\chi}_{n,t,a}^{k,0})^{2},$
subject to $0\leq\eta^{k}_{t,a}\leq 1$, for all $k\in\{0,1,\dots,K-1\}$ and $\sum_{k=0}^{K-1}\eta^{k}_{t,a}=1$, and $\lambda$ is a tuning parameter chosen by cross-validation.
8:  end for
9:  Return:
$\widehat{\theta}^{\text{fed}}_{n}(t,a)=\sum_{k=0}^{K-1}\widehat{\eta}^{k}_{t,a}\widehat{\theta}^{k,0}_{n}(t,a),\text{ and }\widehat{\mathcal{V}}_{t,a}^{\dagger}\text{ for all }(t,a)\in\{0,\epsilon,2\epsilon,\dots,\tau\}\times\{0,1\}.$
$\dagger$: $\widehat{\mathcal{V}}_{t,a}$ is specified in (D.3) in Appendix D.3 by replacing the fixed ${\boldsymbol{\eta}}_{t,a}$ with the estimated value $\widehat{\boldsymbol{\eta}}_{t,a}$.
Algorithm 2 Double/debiased machine learning (DML) algorithm for nuisance function estimations and influence function calculations in Algorithm 1 at a given time point and treatment.
1:  Input: Observed multi-source right-censored data $\mathcal{O}=\{\mathcal{O}_{i}=(\mathbf{X}_{i},A_{i},Y_{i},\Delta_{i},R_{i}),i=1,\dots,n\}=\mathcal{O}^{0}\cup\mathcal{O}^{1}\cup\dots\cup\mathcal{O}^{K-1}$, where $R_{i}\in\{0,1,\dots,K-1\}$ and $\mathcal{O}^{k}$ represents the data for site $R=k$; Given treatment group $A=a$ and a specific time point $t$; The number of disjoint folds into which the data are split, $M$, where $M\in\{2,3,\dots,\lfloor n^{*}/2\rfloor\}$ with $n^{*}=\min\{n,n_{1},\dots,n_{K-1}\}$.
2:  Output: Estimated influence functions for each individual.
3:  Partition $\mathcal{O}^{0}$ into $M$ approximately equal-sized, disjoint validation folds $\mathcal{V}^{0}_{1},\dots,\mathcal{V}^{0}_{M}$, allowing a size difference of at most ±$\pm 1$ between folds.
4:  for $m=1,\dots,M$ do
5:     Define the training set $\mathcal{T}^{0}_{m}=\mathcal{O}^{0}\backslash\mathcal{V}^{0}_{m}$;
6:     Fit nuisance functions $S^{0}$, $G^{0}$, $\pi^{0}$ on $\mathcal{T}^{0}_{m}$, using some methods ensemble from survSuperLearner and SuperLearner;
7:     Predict nuisance functions on $\mathcal{V}^{0}_{m}$ as $\widehat{S}^{0}_{m},\widehat{G}^{0}_{m}$ and $\widehat{\pi}^{0}_{m}$.
8:  end for
9:  Train a model of $S^{0}$ by the entire data of the target site $\mathcal{O}^{0}$, denoted as $S^{0,\text{full}}$, using chosen methods ensemble from survSuperLearner.
10:  for $k=1,\dots,K-1$ do
11:     Partition $\mathcal{O}^{k}$ into $M$ approximately equal-sized, disjoint validation folds $\mathcal{V}^{k}_{1},\dots,\mathcal{V}^{k}_{M}$, allowing a size difference of at most ±$\pm 1$ between folds.
12:     for $m=1,\dots,M$ do
13:        Define the training set $\mathcal{T}^{k}_{m}=\mathcal{O}^{k}\backslash\mathcal{V}^{k}_{m}$;
14:        Fit the density ratio $\omega^{k,0}$ using only covariate data of $\mathcal{T}^{0}_{m}\cup\mathcal{T}^{k}_{m}$, or by just passing through some coarsening level summary statistics;
15:        Fit nuisance functions $G^{k}$, $\pi^{k}$ on $\mathcal{T}^{k}_{m}$, using chosen methods ensembles from survSuperLearner and SuperLearner;
16:        Predict above nuisance functions on $\mathcal{V}^{k}_{m}$ as $\widehat{G}^{k}_{m}$, $\widehat{\omega}^{k,0}_{m}$ and $\widehat{\pi}^{k}_{m}$;
17:        Predict nuisance function $S^{k}$ on $\mathcal{V}^{k}_{m}$ using the pre-trained $S^{0,\text{full}}$ model, and denote the predicted value by $\widehat{S}^{k}_{m}$.
18:     end for
19:     Aggregate all predicted nuisance functions over $M$ folds as $\widehat{S}^{k},\widehat{G}^{k}$, $\widehat{\omega}^{k,0}$ and $\widehat{\pi}^{k}$;
20:  end for
21:  Return: The uncentered EIFs of all individuals in $\mathcal{O}$, by plugging-in their predicted nuisance function values, $\widehat{\varphi}^{k,0}_{t,a}(\mathcal{O};\widehat{S}^{k},\widehat{S}^{0},\widehat{G}^{k},\widehat{\pi}^{k},\widehat{\omega}^{k,0})$, and the centered EIFs $\widehat{\varphi}^{*k,0}_{t,a}(\mathcal{O};\widehat{S}^{k},\widehat{S}^{0},\widehat{G}^{k},\widehat{\pi}^{k},\widehat{\omega}^{k,0})$ by centering the uncentered EIFs to their sample mean, for all $k\in\{0,1,\dots,K-1\}$.
Appendix D Technical Proofs
We adopt the following notation throughout this appendix: (i) $\mathbb{P}_{\infty}$ denotes a general probability limit, and the nuisance functions under $\mathbb{P}_{\infty}$ are denoted with subscript $\infty$, e.g., $S^{0}_{\infty}$; (ii) $\widehat{\mathbb{P}}$ means the corresponding nuisance functions are replaced by their estimates, and $\widehat{\mathbb{P}}$ may converge to a general limit $\mathbb{P}_{\infty}$; (iii) $\mathbb{P}_{n}^{m}[f(\mathcal{O})]=|\mathcal{V}_{m}|^{-1}\sum_{i\in\mathcal{I}_{m}}f(\mathcal{O}_{i})$ to denote the empirical average on the $m$th validation set $\mathcal{V}_{m}$ by cross-fitting, $m=1,\dots,M$.
[…]
D.1 Theory of the CCOD estimator
D.1.1 Proof of Theorem 2.7
Proof.
Recall the estimand $\theta^{0}(t,a)=\mathbb{E}\{S^{0}(t\mid a,\mathbf{X})\mid R=0\}=\mathbb{E}\{\bar{S}(t\mid a,\mathbf{X})\mid R=0\}$ that follows from Assumption 2.5. We write $\bar{\Lambda}(t\mid a,\mathbf{x})=\int_{0}^{t}\frac{\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{D}(u\mid a,\mathbf{x})}$ where $\bar{N}_{\delta}(du\mid a,\mathbf{x})=\mathbb{P}(Y\leq t,\Delta=\delta\mid A=a,\mathbf{X}=\mathbf{x})$ and $\bar{D}(t\mid a,\mathbf{x})=\mathbb{P}(Y\geq t\mid A=a,\mathbf{X}=\mathbf{x})$.
Recall that a mean zero, finite variance function $\varphi^{*0}_{t,a}(\mathcal{O};\mathbb{P})$ is called an influence function of the target estimand (a functional) $\theta^{0}(t,a)=\theta^{0}(t,a;\mathbb{P})$ at $\mathbb{P}$ if, for any one-dimensional regular parametric submodel $\{\mathbb{P}_{\epsilon}:\epsilon\in[0,1)\}$ through $\mathbb{P}\equiv\mathbb{P}_{0}$,
$\left.\frac{\partial}{\partial\epsilon}\theta^{0}(t,a;\mathbb{P}_{\epsilon})\right|_{\epsilon=0}=\mathbb{E}_{\mathbb{P}}[\varphi^{*0}_{t,a}(\mathcal{O};\mathbb{P})\dot{\ell}(\mathcal{O})],$
where ˙$\dot{\ell}(\mathcal{O})$ is the score function of the submodel at $\epsilon=0$ (i.e., typically, ˙$\dot{\ell}(\mathcal{O})={\partial}\log{\{p_{\epsilon}(\mathcal{O})\}}/{\partial\epsilon}~{}|_{\epsilon=0}$), where $p_{\epsilon}(\cdot)$ denotes the probability density (likelihood) function under submodel $\mathbb{P}_{\epsilon}$ [4]. Thus, to find the EIF, we begin by writing the equation
00 0 $=\frac{\partial}{\partial\epsilon}\theta^{0}(t,a;\mathbb{P}_{\epsilon})\bigg{|}_{\epsilon=0}=\frac{\partial}{\partial\epsilon}\mathbb{E}_{\mathbb{P}_{\epsilon}}\{\bar{S}_{\epsilon}(t\mid a,\mathbf{X})\mid R=0\}\bigg{|}_{\epsilon=0}$
$=\mathbb{E}\{[\bar{S}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}_{\mathbf{X}\mid R=0}\mid R=0\}+\mathbb{E}\left\{\int\frac{\partial}{\partial\epsilon}\bar{S}_{\epsilon}(t\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}\mu(d\mathbf{x})~{}\bigg{|}~{}R=0\right\},$ (3)
where $\mu(\cdot)$ denotes the distribution of $\mathbf{X}$ induced by $\mathbb{P}$ and, for any sets of variables $V$ and $W$, ˙$\dot{\ell}_{V\mid W}$ denotes the conditional score function of $V$ given $W$, i.e., typically $\partial\log{\{p_{\epsilon}(V\mid W)\}}/{\partial\epsilon}~{}|_{\epsilon=0}$—note that such scores always satisfy $\mathbb{E}_{\mathbb{P}}(\dot{\ell}_{V\mid W}\mid W)=0$ [4].
[…]
$\alpha\mapsto\bar{S}(t;H)\int_{0}^{t}\frac{\bar{S}(u-;H)}{\bar{S}(u;H)}\alpha(du)$
at $H$.
Thus, by letting $H(t)=\bar{\Lambda}_{\epsilon}(t\mid a,\mathbf{x})$ and the chain rule, the integrand in the second term becomes
[…]
Furthermore,
$\frac{\partial}{\partial\epsilon}\bar{\Lambda}_{\epsilon}(du\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\dfrac{\frac{\partial}{\partial\epsilon}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})\mid_{\epsilon=0}}{\bar{D}(u\mid a,\mathbf{x})}-\dfrac{\frac{\partial}{\partial\epsilon}\bar{D}_{\epsilon}(u\mid a,\mathbf{x})\mid_{\epsilon=0}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})}{\bar{D}(u\mid a,\mathbf{x})^{2}}.$
In addition,
$\frac{\partial}{\partial\epsilon}\bar{N}_{1,\epsilon}(du\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\frac{\partial}{\partial\epsilon}\mathbb{P}_{\epsilon}(Y\leq u,\Delta=1\mid A=a,\mathbf{X}=\mathbf{x})\bigg{|}_{\epsilon=0}$
$=\frac{\partial}{\partial\epsilon}\iint\mathbb{I}(y\leq u,\delta=1)\mathbb{P}_{\epsilon}(dy,d\delta\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$
$=\iint\mathbb{I}(y\leq u,\delta=1)\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x})$
$=\int_{\delta}\mathbb{I}(\delta=1)\dot{\ell}(u,\delta\mid a,\mathbf{x},k)\mathbb{P}(du,d\delta\mid a,\mathbf{x}),$
and
$\frac{\partial}{\partial\epsilon}\bar{D}_{\epsilon}(u\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$ $=\frac{\partial}{\partial\epsilon}\mathbb{P}_{\epsilon}(Y\geq u\mid A=a,\mathbf{X}=\mathbf{x})\bigg{|}_{\epsilon=0}$
$=\frac{\partial}{\partial\epsilon}\iint\mathbb{I}(y\geq u)\mathbb{P}_{\epsilon}(dy,d\delta\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}$
$=\iint\mathbb{I}(y\leq u)\dot{\ell}(y,\delta\mid a,\mathbf{x})\mathbb{P}(dy,d\delta\mid a,\mathbf{x}).$
Therefore, plugging-in the above expressions,
[…]
$=\iiint-\mathbb{I}(y\leq t,\delta=1)\frac{\bar{S}(t\mid a,\mathbf{x})\bar{S}(y-\mid a,\mathbf{x})}{\bar{S}(y\mid a,\mathbf{x})\bar{D}(y\mid\mathbf{x})}\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x})\mu(d\mathbf{x})$
$\quad+\iiiint\mathbb{I}(u\leq t,u\leq y)\frac{\bar{S}(t\mid a,\mathbf{x})\bar{S}(u-\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid\mathbf{x})}$
×˙$\quad\qquad\times\dot{\ell}(y,\delta\mid a,\mathbf{x})\mathbb{P}(dy,d\delta\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})\mu(d\mathbf{x})$
$=\iiint-\mathbb{I}(y\leq t,\delta=1)\frac{\bar{S}(t\mid a,\mathbf{x})\bar{S}(y-\mid a,\mathbf{x})}{\bar{S}(y\mid a,\mathbf{x})\bar{D}(y\mid\mathbf{x})}\dot{\ell}(y,\delta\mid a,\mathbf{x})\mathbb{P}(dy,d\delta\mid a,\mathbf{x})\mu(d\mathbf{x})$
$\quad+\iiint\bar{S}(t\mid a,\mathbf{x})\int_{0}^{t\wedge y}\frac{\bar{S}(u-\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid\mathbf{x})^{2}}\bar{N}_{1}(du\mid a,\mathbf{x})$
×˙$\quad\qquad\times\dot{\ell}(y,\delta\mid a,\mathbf{x})\mathbb{P}(dy,d\delta\mid a,\mathbf{x})\mu(d\mathbf{x})$
$=\mathbb{E}\left[\bar{S}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\left\{\bar{H}(t\wedge Y,a,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)\bar{S}(Y-\mid a,\mathbf{X})}{\bar{S}(Y\mid a,\mathbf{X})\bar{D}(Y\mid a,\mathbf{X})}\right\}\dot{\ell}(Y,\Delta\mid a,\mathbf{X})\right],$
where
$\bar{H}(t,a,\mathbf{x})=\int_{0}^{t}\frac{\bar{S}(u-\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid a,\mathbf{x})^{2}}.$
Now, we note that
$\mathbb{E}\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)\bar{S}(Y-\mid A,\mathbf{X})}{\bar{S}(Y\mid A,\mathbf{X})\bar{D}(Y\mid A,\mathbf{X})}~{}\bigg{|}~{}A=a,\mathbf{X}=\mathbf{x}\right]=\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{x})\bar{N}_{1}(dy\mid a,\mathbf{x})}{\bar{S}(y\mid a,\mathbf{x})\bar{D}(y\mid a,\mathbf{x})},$
and
$\mathbb{E}\{\bar{H}(t\wedge Y,A,\mathbf{X})\mid A=a,\mathbf{X}=\mathbf{x}\}$
$=\iint^{t}\mathbb{I}(u\leq y)\frac{\bar{S}(u-\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid a,\mathbf{x})^{2}}\mathbb{P}(dy\mid a,\mathbf{x})$
$=\int_{0}^{t}\mathbb{P}(Y\geq u\mid A=a,\mathbf{X}=\mathbf{x})\frac{\bar{S}(u-\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid a,\mathbf{x})^{2}}\mathbb{P}(dy\mid a,\mathbf{x})$
$=\int_{0}^{t}\frac{\bar{S}(u-\mid a,\mathbf{x})\bar{N}_{1}(du\mid a,\mathbf{x})}{\bar{S}(u\mid a,\mathbf{x})\bar{D}(u\mid a,\mathbf{x})}.$
Therefore,
$\mathbb{E}\left[\bar{H}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)\bar{S}(Y-\mid A,\mathbf{X})}{\bar{S}(Y\mid A,\mathbf{X})\bar{D}(Y\mid A,\mathbf{X})}~{}\bigg{|}~{}A,\mathbf{X}\right]=0$
almost surely. By the fact that score functions have (conditional) mean zero, and by the tower property of expectation, the above calculations implies that
[…]
$=\mathbb{E}\left[\bar{S}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\left\{\bar{H}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)\bar{S}(Y-\mid A,\mathbf{X})}{\bar{S}(Y\mid A,\mathbf{X})\bar{D}(Y\mid A,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})\right].$
Combining these results with the facts that $\bar{N}_{1}(du\mid a,\mathbf{x})/\bar{D}(u\mid a,\mathbf{x})=\bar{\Lambda}(du\mid a,\mathbf{x})$ and $\bar{D}(u\mid a,\mathbf{x})=\bar{S}(u-\mid x)\bar{G}(u\mid a,\mathbf{x})$, we can rewrite (D.1.1) at the beginning as follows:
$\frac{\partial}{\partial\epsilon}\theta^{0}(t,a;\mathbb{P}_{\epsilon})\bigg{|}_{\epsilon=0}$
$=\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}[\bar{S}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}(\mathcal{O})-\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\mathbb{E}\bigg{\{}\bar{S}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}$
×$\quad\quad\times\left\{\frac{\mathbb{I}(Y\leq t,\Delta=1)}{\bar{S}(y\mid\mathbf{X})\bar{G}(y\mid a,\mathbf{X})}-\int_{0}^{t\wedge y}\frac{\bar{\Lambda}(du\mid a,\mathbf{X})}{\bar{S}(u\mid\mathbf{X})\bar{G}(u\mid a,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})~{}\bigg{|}~{}\mathbf{X}\bigg{\}}\bigg{]}$
$=\mathbb{E}\left[\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}[\bar{S}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}(\mathcal{O})\right]-\mathbb{E}\bigg{[}\frac{\mathbb{P}(R=0\mid\mathbf{X})}{\mathbb{P}(R=0)}\bar{S}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}$
×$\quad\quad\times\left\{\frac{\mathbb{I}(Y\leq t,\Delta=1)}{\bar{S}(y\mid\mathbf{X})\bar{G}(y\mid a,\mathbf{X})}-\int_{0}^{t\wedge y}\frac{\bar{\Lambda}(du\mid a,\mathbf{X})}{\bar{S}(u\mid\mathbf{X})\bar{G}(u\mid a,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})\bigg{]}.$
Therefore, the (uncentered) EIF of the target estimand $\theta^{0}(t,a)$ at $\mathbb{P}$ is found as
$\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\bar{S},\bar{G},\bar{\pi})$ $=\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P})$
$=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\bar{S}(t\mid a,\mathbf{X})-\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\bar{S}(t\mid a,\mathbf{X})$
×$\quad\times\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)}{\bar{S}(Y\mid a,\mathbf{X})\bar{G}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\bar{\Lambda}(du\mid a,\mathbf{X})}{\bar{S}(u\mid a,\mathbf{X})\bar{G}(u\mid a,\mathbf{X})}\right],$
where $q^{0}(\mathbf{X})=\mathbb{P}(R=0\mid\mathbf{X})$ is the target site propensity. ∎
D.1.2 Regularity conditions for the CCOD estimator
The proof of Theorem 2.8 requires the following regularity conditions with $\bar{\pi}_{\infty}=\bar{\pi}$, $q^{0}_{\infty}=q^{0}$, $\bar{G}_{\infty}=\bar{G}$ and $\bar{S}_{\infty}=\bar{S}$.
Condition D.1.
There exists $\bar{\pi}_{\infty}$, $\bar{G}_{\infty}$, $\bar{\Lambda}_{\infty}$ and $\bar{S}_{\infty}$ such that
  1. (a)
    $\max_{m}\mathbb{P}\left[\frac{1}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}-\frac{1}{\bar{\pi}_{\infty}(a\mid\mathbf{X})}\right]^{2}\to_{p}0$,
  2. (b)
    $\max_{m}\mathbb{P}\left[{\widehat{q}^{0}_{m}(\mathbf{X})}-{q^{0}_{\infty}(\mathbf{X})}\right]^{2}\to_{p}0$,
  3. (c)
    $\max_{m}\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{1}{\widehat{\bar{G}}_{m}(u\mid a,\mathbf{X})}-\frac{1}{\bar{G}_{\infty}(u\mid a,\mathbf{X})}\right|\right]^{2}\to_{p}0$, and
  4. (d)
    $\max_{m}\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})}{\widehat{\bar{S}}_{m}(u\mid a,\mathbf{X})}-\frac{\bar{S}_{\infty}(t\mid a,\mathbf{X})}{\bar{S}_{\infty}(u\mid a,\mathbf{X})}\right|\right]^{2}\to_{p}0$.
Condition D.2.
There exists $\eta\in(0,\infty)$ such that for $\mathbb{P}$-almost all $\mathbf{x}$, $\widehat{\bar{\pi}}_{m}(a\mid\mathbf{x})\geq 1/\eta$, $\bar{\pi}_{\infty}(a\mid\mathbf{x})\geq 1/\eta$, $\widehat{q}^{0}_{m}(\mathbf{x})\geq 1/\eta$, $q^{0}_{\infty}(\mathbf{x})\geq 1/\eta$, $\widehat{\bar{G}}_{m}(t\mid a,\mathbf{x})\geq 1/\eta$, and $\bar{G}_{\infty}(t\mid a,\mathbf{x})\geq 1/\eta$ with probability tending to 1.
Condition D.3.
Define
$\bar{r}_{n,t,a,1}$ $=\max_{m}\mathbb{P}|\{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})-\bar{\pi}(a\mid\mathbf{X})\}\cdot\{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}|,$
$\bar{r}_{n,t,a,2}$ $=\max_{m}\mathbb{P}|\{\widehat{q}^{0}_{m}(\mathbf{X})-q^{0}(\mathbf{X})\}\cdot\{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}|,\text{ and}$
$\bar{r}_{n,t,a,3}$ $=\max_{m}\mathbb{P}\bigg{|}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\left\{\frac{\bar{G}(u\mid a,\mathbf{X})}{\widehat{\bar{G}}_{m}(u\mid a,\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{|}.$
Then, it holds that all $\bar{r}_{n,t,a,1}=o_{p}(n^{-1/2})$, $\bar{r}_{n,t,a,2}=o_{p}(n^{-1/2})$ and $\bar{r}_{n,t,a,3}=o_{p}(n^{-1/2})$.
Condition D.4.
$\max_{m}\mathbb{P}\left[\sup_{u\in[0,t]}\sup_{v\in[0,u]}\left|\frac{\widehat{\bar{S}}_{m}(u\mid a,\mathbf{X})}{\widehat{\bar{S}}_{m}(v\mid a,\mathbf{X})}-\frac{\bar{S}_{\infty}(u\mid a,\mathbf{X})}{\bar{S}_{\infty}(v\mid a,\mathbf{X})}\right|\right]^{2}\to_{p}0.$
Condition D.5.
It holds that all $\sup_{u\in[0,t]}\bar{r}_{n,u,a,1}=o_{p}(n^{-1/2})$, $\sup_{u\in[0,t]}\bar{r}_{n,u,a,2}=o_{p}(n^{-1/2})$, and $\sup_{u\in[0,t]}\bar{r}_{n,u,a,3}=o_{p}(n^{-1/2})$.
D.1.3 Lemmata for the CCOD estimator
To establish the RAL related results of the CCOD estimator, we first present a number of lemmata in this section. We assume that the corresponding nuisance functions converge to some general limits defined by $\mathbb{P}_{\infty}$. We start by expressing the difference $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)-\theta^{0}(t,a)$ as
$\mathbb{P}_{n}[\widehat{\varphi}^{\text{CCOD}}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})]-\theta^{0}(t,a)$
$=\frac{1}{n}\sum_{m=1}^{M}\sum_{i\in\mathcal{I}_{m}}\widehat{\varphi}^{\text{CCOD}}_{t,a}(\mathcal{O}_{i})-\theta^{0}(t,a)$
$=\mathbb{P}_{n}[\varphi^{\text{CCOD}}_{\infty,t,a}]-\theta^{0}(t,a)+\frac{1}{n}\sum_{m=1}^{M}\sum_{i\in\mathcal{I}_{m}}\widehat{\varphi}^{\text{CCOD}}_{t,a}(\mathcal{O}_{i})-\mathbb{P}_{n}[\varphi^{\text{CCOD}}_{t,a}]$
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{n}\sum_{m=1}^{M}\sum_{i\in\mathcal{L}^{m}}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}(\mathcal{O}_{i})-\varphi^{\text{CCOD}}_{\infty,t,a}(\mathcal{O}_{i})\right]$
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\frac{1}{n_{m}}\sum_{i\in\mathcal{L}^{m}}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}(\mathcal{O}_{i})-\varphi^{\text{CCOD}}_{\infty,t,a}(\mathcal{O}_{i})\right]$
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}(\mathbb{P}_{n}^{m}-\mathbb{P})\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$
$=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\theta^{0}(t,a)\right].$ (4)
Next, in the following Lemma D.1, we establish an $L_{2}(\mathbb{P})$ norm distance (bound) between the estimated IF and the limiting IF in terms of discrepancies on the nuisance parameters, we consider some decompositions.
Lemma D.1.
Under Condition D.2, there exists a universal constant $C=C(\eta)$ such that for each $m$, $n$, $t$, and $a$,
$\mathbb{P}[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}]^{2}\leq C(\eta)\sum_{j=1}^{6}\bar{A}_{j,n,m,t,a},$
where
$\bar{A}_{1,n,m,t,a}$ $=\mathbb{P}\left[\frac{1}{\mathbb{P}_{n}^{m}(R=0)}-\frac{1}{\mathbb{P}(R=0)}\right]^{2},$
$\bar{A}_{2,n,m,t,a}$ $=\mathbb{P}\left[\widehat{q}^{0}_{m}(\mathbf{X})-q^{0}_{\infty}(\mathbf{X})\right]^{2},$
$\bar{A}_{3,n,m,t,a}$ $=\mathbb{P}\left[\frac{1}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}-\frac{1}{\bar{\pi}_{\infty}(a\mid\mathbf{X})}\right]^{2},$
$\bar{A}_{4,n,m,t,a}$ $=\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{1}{\widehat{\bar{G}}_{m}(u\mid a,\mathbf{X})}-\frac{1}{\bar{G}_{\infty}(u\mid a,\mathbf{X})}\right|\right]^{2},$
$\bar{A}_{5,n,m,t,a}$ $=\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})}{\widehat{\bar{S}}_{m}(u\mid a,\mathbf{X})}-\frac{\bar{S}_{\infty}(t\mid a,\mathbf{X})}{\bar{S}_{\infty}(u\mid a,\mathbf{X})}\right|\right]^{2}.$
Proof.
We first denote
$\bar{B}(\mathcal{V}_{m})$ $=\frac{\mathbb{I}(A=a)}{\bar{\pi}(a\mid\mathbf{X})}\bar{S}(t\mid a,\mathbf{X})\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)}{\bar{S}(Y\mid a,\mathbf{X})\bar{G}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\bar{\Lambda}(du\mid a,\mathbf{X})}{\bar{S}(u\mid a,\mathbf{X})\bar{G}(u\mid a,\mathbf{X})}\right].$
Then, we have the following decomposition:
$\widehat{\varphi}^{\text{CCOD}}_{t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}=\sum_{j=1}^{4}\bar{U}_{j,n,m,t,a},$
where
$\bar{U}_{1,n,m,t,a}$ $=\left[\frac{\mathbb{I}(R=0)}{\mathbb{P}_{n}^{m}(R=0)}-\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\right]\widehat{\bar{S}}_{m}(t\mid a,\mathbf{x}),$
$\bar{U}_{2,n,m,t,a}$ $=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\left[\widehat{\bar{S}}_{m}(t\mid a,\mathbf{x})-\bar{S}_{\infty}(t\mid a,\mathbf{x})\right],$
$\bar{U}_{3,n,m,t,a}$ $=\left[\frac{\widehat{q}^{0}_{m}(\mathbf{X})}{\mathbb{P}_{n}^{m}(R=0)}-\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\right]\widehat{\bar{B}}_{m}(\mathcal{V}_{m}),$
$\bar{U}_{4,n,m,t,a}$ $=\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\left[\widehat{\bar{B}}_{m}(\mathcal{V}_{m})-\bar{B}_{\infty}(\mathcal{V}_{m})\right].$
Note that the expression of $\widehat{\bar{B}}_{m}(\mathcal{V}_{m})-\bar{B}_{\infty}(\mathcal{V}_{m})$ can be found in Lemma 3 of Westling et al. [48], while we only need to replace corresponding nuisance functions with the global version here, thus the detail is omitted. Then, by the triangle inequality, we have $\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]^{2}\leq\left[\sum_{j=1}^{4}\{\mathbb{P}\{(\bar{U}_{j,n,m,t,a})^{2}\}\}^{1/2}\right]^{2}$. Therefore, under Assumption 2.3 and Condition D.2, there exists a universal constant $C=C(\eta)$ such that the result in the statement holds. Thus, the proof is completed. ∎
Furthermore, we investigate conditions that make the empirical process term $\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]$ to be $o_{p}(n^{-1/2})$. This requires some preliminaries from the empirical process theory [43]. We first introduce the following notation and Lemma D.2. Then, we consider results for uniform convergence over $t\in[0,\tau]$.
[…]
Lemma D.2 (Lemma 4 in Westling et al. [48]).
Let =$\mathcal{F}=\{x\mapsto f_{t}(x):t\in[0,\tau]\}$ be a class of functions on a sample space $\mathcal{X}$ such that $f_{s}(x)\leq f_{t}(x)$ for all $0\leq s\leq t\leq\tau$ and $x\in\mathcal{X}$, and such that an envelope $F$ for $\mathcal{F}$ satisfies $\|F\|_{\mathbb{P},2}<\infty$. Then $N_{[]}(\varepsilon\|F\|_{\mathbb{P},2},\mathcal{F},L_{2}(\mathbb{P}))\leq\mathbb{P}(f_{\tau}-f)^{2}/(\varepsilon^{2}\|F\|_{\mathbb{P},2}^{2})\leq 4\varepsilon^{2}$ for all $(0,1]$. If $\mathcal{F}$ is uniformly bounded by a constant $C$, then $N(\varepsilon C,\mathcal{F},L_{2}(\mathbb{Q}))\leq 1/\varepsilon^{2}$ for each probability distribution $\mathbb{Q}$ on $\mathcal{X}$.
Then, we introduce the following Lemma D.3. For simplicity, $p^{0}$ and $q^{0}$ correspond to, respectively, $\mathbb{P}(R=0)$ and $q^{0}(\mathbf{X})$, and $S,\pi,G$ correspond to, respectively $\bar{S}$, $\bar{\pi}$ and $\bar{G}$ functions in Lemma D.3.
Lemma D.3.
Let $S,\pi,q^{0},p^{0}$, and $G$ be fixed, where $t\mapsto S(t\mid a,\mathbf{x})$ is assumed to be non-increasing for each $(a,\mathbf{x})$, and where $S(t_{0}\mid a,\mathbf{x})\geq 1/\eta$, $G(t_{0}\mid a,\mathbf{x})\geq 1/\eta$, $\pi(a_{0}\mid\mathbf{x})\geq 1/\eta$, and $p^{0}\geq 1/\eta$, for some $\eta\in(0,\infty)$. Then the class of influence functions $\mathcal{F}_{S,\pi,p^{0},q^{0},G,t_{0},a_{0}}=\{\varphi_{S,\pi,p^{0},q^{0},G,t_{0},a_{0}}:t\in[0,t_{0}]\}$ satisfies
$\sup_{\mathbb{Q}}N(\varepsilon\|F\|_{\mathbb{Q},2},\mathcal{F}_{S,\pi,p^{0},q^{0},G,t_{0},a_{0}},L_{2}(\mathbb{Q}))\leq 32/\varepsilon^{10},$
for any $\varepsilon\in(0,1]$ where $F=\eta(1+2\eta^{2})$ is an envelope of $\mathcal{F}_{S,\pi,p^{0},q^{0},G,t_{0},a_{0}}$, and the supremum is taken over all distributions $\mathbb{Q}$ on the sample space of the observed data.
Proof.
The class $\mathcal{F}_{S,\pi,p^{0},q^{0},G,t_{0},a_{0}}$ is uniformly bounded by $\eta(1+2\eta^{2})$ because of the assumed upper bounds of $1/p^{0},1/\pi$ and $1/G$. Therefore, the envelop function can be taken as $F=\eta(1+2\eta^{2})$.
Define the following two functions $f_{t}$ and $h_{t}$ pointwise as
$f_{t}(\mathbf{x},a_{0},\delta,y)$ $=\frac{\mathbb{I}(a=a_{0},y\leq t,\delta=1)q^{0}(\mathbf{x})S(t\mid a_{0},\mathbf{x})}{p^{0}\pi(a_{0}\mid\mathbf{x})S(y\mid a_{0},\mathbf{x})G(y\mid a_{0},\mathbf{x})},$
$h_{t}(\mathbf{x},a_{0},y)$ $=\int\frac{\mathbb{I}(a=a_{0},u\leq y,u\leq t)q^{0}(\mathbf{x})S(t\mid a_{0},\mathbf{x})}{p^{0}\pi(a_{0}\mid\mathbf{x})S(u\mid a_{0},\mathbf{x})G(u\mid a_{0},\mathbf{x})}\Lambda(du\mid a_{0},\mathbf{x}).$
Consider classes $\mathcal{F}^{1}_{S,p^{0},q^{0},t_{0},a_{0}}=\{(\mathbf{x},r)\mapsto\mathbb{I}(r=0)S(t\mid a_{0},\mathbf{x})/p^{0}:t\in[0,t_{0}]\}$, $\mathcal{F}^{2}_{S,p^{0},q^{0},G,t_{0},a_{0}}=\{(\mathbf{x},a_{0},\delta,y)\mapsto f_{t}(\mathbf{x},a_{0},\delta,y):t\in[0,t_{0}]\}$, and $\mathcal{F}^{3}_{S,p^{0},q^{0},G,t,a_{0}}=\{(\mathbf{x},a_{0},y)\mapsto h_{t}(\mathbf{x},a,y):t\in[0,t_{0}]\}$. We can then write
$\mathcal{F}_{S,\pi,p^{0},q^{0},G,t_{0},a_{0}}\subseteq\{f_{1}-f_{2}+f_{3}:f_{1}\in\mathcal{F}^{1}_{S,p^{0},q^{0},t_{0},a_{0}},f_{2}\in\mathcal{F}^{2}_{S,p^{0},q^{0},G,t_{0},a_{0}},f_{3}\in\mathcal{F}^{3}_{S,p^{0},q^{0},G,t_{0},a_{0}}\}.$
Since $(\mathbf{x},r)\mapsto\mathbb{I}(r=0)S(t\mid a_{0},\mathbf{x})/p^{0}$ is non-increasing over $t\in[0,t_{0}]$ for all $(\mathbf{x},r)$ and uniformly bounded by $1/\eta$, Lemma D.2 implies that $\sup_{\mathbb{Q}}(\varepsilon,\mathcal{F}^{1}_{S,p^{0},q^{0},t_{0},a_{0}},L_{2}(\mathbb{Q}))\leq 2\varepsilon^{2}$.
The $\mathcal{F}^{2}_{S,p^{0},q^{0},G,t_{0},a_{0}}$ is contained in the product of the classes $\{y\mapsto\mathbb{I}(y\leq t):t\in[0,t_{0}]\}$, $\{\mathbf{x}\mapsto S(t\mid a_{0},\mathbf{x}):t\in[0,t_{0}]\}$, and the singleton class $\{(\mathbf{x},a_{0},\delta,y)\mapsto\mathbb{I}(a=a_{0},\delta=1,y\leq t)q^{0}(\mathbf{x})/[p^{0}\pi(a_{0}\mid\mathbf{x})S(y\mid a_{0},\mathbf{x})G(y\mid a_{0},\mathbf{x})]\}.$ The first two classes both have covering number $\sup_{\mathbb{Q}}N(\varepsilon,\cdot,L_{2}(\mathbb{Q}))\leq 2\varepsilon^{2}$ by Lemma D.2. The third class has uniform covering number 1 for all $\varepsilon$ because it can be covered with a single ball of any positive radius. In addition, $\mathcal{F}^{2}_{S,p^{0},q^{0},G,t,a}$ is uniformly bounded by $\eta^{3}$. Therefore, based on Lemma 5.1 in Van der Laan [44], $\sup_{\mathbb{Q}}N(\varepsilon\eta^{3},\mathcal{F}^{2}_{S,p^{0},q^{0},G,t,a},L_{2}(\mathbb{Q}))\leq 4/\varepsilon^{4}$.
Furthermore, based on Lemma 5.1 in Van der Laan [44] again, we can rewrite $\mathcal{F}^{3}_{S,p^{0},q^{0},G,t_{0},a_{0}}=\{(\mathbf{x},a_{0},y)\mapsto\int m_{t}(u,\mathbf{x},a_{0},y)\mu^{*}(du):t\in[0,t_{0}]\}$, where
$m_{t}(u,\mathbf{x},a_{0},y)=\frac{\mathbb{I}(a=a,u\leq y,u\leq t)q^{0}(\mathbf{x})S(t\mid a_{0},\mathbf{x})}{p^{0}\pi(a_{0}\mid\mathbf{x})S(u\mid a_{0},\mathbf{x})G(u\mid a_{0},\mathbf{x})}\lambda^{*}(u\mid a_{0},\mathbf{x}),$
with $\mu^{*}(du)$ is a dominating measure for $\mathbb{P}$-almost all $\mathbf{x}$ and $\lambda^{*}(du\mid a_{0},\mathbf{x})$ the Radon-Nikodym derivative of $\Lambda(\cdot\mid a_{0},\mathbf{x})$ with respect to $\mu^{*}$. Then, the class $\mathcal{M}_{t}=\{m_{t}:t\in[0,t_{0}]\}$ is contained in the product of the singleton class $\{(\mathbf{x},a_{0},y,u)\mapsto\mathbb{I}(a=a_{0},u\leq y)q^{0}(\mathbf{x})\lambda^{*}(u\mid a_{0},\mathbf{x})/[p^{0}\pi(a_{0}\mid\mathbf{x})S(u\mid a_{0},\mathbf{x})G(u\mid a_{0},\mathbf{x})]\}$ and the class $\{u\mapsto\mathbb{I}(u\leq t):t\in[0,t_{0}]\}$ and $\{\mathbf{x}\mapsto S(t\mid a_{0},\mathbf{x}):t\in[0,t]\}$, which as discussed above both have $L_{2}(\mathbb{Q})$ covering number bounded by $2/\varepsilon^{2}$ for any probability measure $\mathbb{Q}$. Therefore, $\sup_{\mathbb{Q}}N(\varepsilon\eta^{3},\mathcal{M}_{t},L_{2}(\mathbb{Q}))\leq 4/\varepsilon^{4}$. By Jensen’s inequality, $\|h_{t}-h_{s}\|_{L_{2}(\mathbb{Q})}\leq\|m_{t}-m_{s}\|_{L_{2}(\mu^{*}\times\mathbb{Q})}$ for any $\mathbb{Q}$, which implies that $\sup_{\mathbb{Q}}N(\varepsilon\eta^{3},\mathcal{F}^{3}_{S,p^{0},q^{0},G,t,a},L_{2}(\mathbb{Q}))\leq\sup_{\mathbb{Q}}N(\varepsilon\eta^{3},\mathcal{M}_{t},L_{2}(\mu^{*}\times\mathbb{Q}))\leq 4/\varepsilon^{4}$ for all $\mathbb{Q}$.
Therefore, we have shown that the three classes have covering numbers bounded by $2/\varepsilon^{2}$, $4/\varepsilon^{4}$ and $4/\varepsilon^{4}$, respectively. Therefore, by Lemma 5.1 in Van der Laan [44],
$\sup_{\mathbb{Q}}N(\varepsilon\eta(1+2\eta^{2}),\mathcal{F}_{S,\pi,p^{0},q^{0},G,t_{0},a_{0}},L_{2}(\mathbb{Q}))\leq 32/\varepsilon^{10}.$
Lemma D.4.
If Conditions D.1D.2 hold, $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}^{1/2}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a_{0}}\right]=o_{p}(n^{-1/2})$. If Conditions D.4 holds as well,
$\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a_{0}}\right]\right|=o_{p}(n^{-1/2}).$
Proof.
We follow notation in Lemma D.1. First, we note that
$\frac{Mn_{m}^{1/2}}{n}$ $\leq\frac{M(|n_{m}-n/M|+n/M)^{1/2}}{n}$
$\leq\frac{M|n_{m}-n/M|^{1/2}+M|n/M|^{1/2}}{n}$
$\leq\left(\frac{M}{n}\right)^{1/2}+\frac{M}{n},$
for all $m$ since $|n_{m}-n/M|\leq 1$ by assumption on $n_{m}$. Then, we have that
$\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|$
$\leq O(n^{-1/2})\frac{1}{M}\sum_{m=1}^{M}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|,$
since $K=O(1)$.
Therefore, for the pointwise claim, we turn to show $M^{-1}\sum_{m=1}^{M}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|=o_{p}(1)$. Using conditional argument, we write
$\mathbb{E}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|=\mathbb{E}\left[\mathbb{E}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|\mid\mathcal{T}_{m}\right],$
where $\mathcal{T}_{m}=\mathcal{O}\backslash\mathcal{V}_{m}$ is the $m$th training set. Note that the randomness in the inner expectation of the right-hand-side above, by conditioning on the training set, is only induced from $\mathbb{G}_{n}^{m}$ by averaging over the observations on the validation set. Therefore,
[…]
Defining $\mathcal{F}^{\text{CCOD}}_{n,m,t,a}$ as the singleton class of functions $\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a}$, we further have
$\mathbb{P}\left|\mathbb{G}_{n}^{m}(\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a})\right|=\mathbb{P}\left[\sup_{f\in\mathcal{F}^{\text{CCOD}}_{n,m,t,a}}\left|\mathbb{G}_{n}^{m}(f)\right|\right].$
By Theorem 2.1.14 in van der Vaart and Wellner [46], the covering number of $\mathcal{F}^{\text{CCOD}}_{n,m,t,a}$ is 1 for all $\varepsilon$, so the uniform entropy integral $J(1,\mathcal{F}^{\text{CCOD}}_{n,m,t,a})$ is 1 relative to the natural envelope $|\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}|$. Therefore, there is a universal constant $C^{\prime}$ such that
$\mathbb{P}\left[\sup_{f\in\mathcal{F}^{\text{CCOD}}_{n,m,t,a}}\left|\mathbb{G}_{n}^{m}(f)\right|\right]\leq C^{\prime}\left\{\mathbb{P}(\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a})^{2}\right\}^{1/2}\leq C^{\prime\prime}\sum_{j=1}^{6}\bar{A}_{j,n,m,t,a},$
following definition of $\bar{A}_{j,n,m,t,a}$ terms in Lemma D.8, so that $M^{-1}\sum_{m=1}^{M}\mathbb{E}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|$ is bounded up to $C^{\prime\prime\prime}\sum_{j=1}^{6}\mathbb{E}[\max_{m}(\bar{A}_{j,n,m,t,a})]$ for some constant $C^{\prime\prime\prime}$. It is straightforward that by Conditions D.6 and D.7, this upper bound tends to zero, so $M^{-1}\sum_{m=1}^{M}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{\infty,t,a}\right]\right|=o_{p}(1)$.
Next, we show the uniform statement. The basic argument is the same. We first write
$\mathbb{E}\left[\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left\{\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}\right\}\right|~{}\bigg{|}~{}\mathcal{T}_{m}\right]$ $=\mathbb{E}\left[\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}\right]\right|\right]$
$=\mathbb{E}\left[\sup_{g\in\mathcal{G}_{n,m,t,a}}\left|\mathbb{G}_{n}^{m}g\right|\right],$
where $\mathcal{G}_{n,m,t,a}=\{\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}:u\in[0,t]\}$. When conditioning on $\mathcal{T}_{m}$, the functions $\widehat{\bar{S}}_{m}$, $\widehat{\bar{G}}_{m}$, $\widehat{\bar{\pi}}_{m}$ and $\widehat{q}^{0}_{m}$ are fixed, so Lemma D.3 implies that
$\log\sup_{\mathbb{Q}}N(\varepsilon\|\bar{G}_{n,m,t,a}\|_{\mathbb{Q},2},\mathcal{G}_{n,m,t,a},L_{2}(\mathbb{Q}))\leq\widetilde{C}\log\varepsilon^{-1},$
for some constant $\widetilde{C}$ not depending on $n,m,$ or $\varepsilon$, and where $\bar{G}_{n,m,t,a}:=\sup_{u\in[0,t]}|\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}|$ is the natural envelope function for $\mathcal{G}_{n,m,t,a}$. As a result, the uniform entropy integral
[…]
is bounded by a constant not depending on $n$ or $m$. By Theorem 2.1.14 in Van der Vaart and Wellner [46], there is therefore a constant $\bar{C}$ not not depending on $n$ or $m$ such that
$\mathbb{E}\left[\sup_{g\in\mathcal{G}_{n,m,t,a}}\left|\mathbb{G}_{n}^{m}g\right|\right]$ $\leq\bar{C}\mathbb{P}\left[\sup_{u\in[0,t]}\{\varphi^{\text{CCOD}}_{n,m,u,a}(\mathcal{O})-\varphi^{\text{CCOD}}_{\infty,u,a}(\mathcal{O})\}^{2}\right]^{2}$
$\leq\bar{C}C(\eta)\sum_{j=1}^{6}\bar{A}_{j,n,m,t,a},$
where the second inequality follows notation and results of Lemma D.1. We therefore have that
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{E}\left[\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,u,a}-\varphi^{\text{CCOD}}_{\infty,u,a}\right]\right|\right]\leq\bar{C}C(\eta)\sum_{j=1}^{6}\max_{m}\mathbb{E}[\bar{A}_{j,n,m,t,a}].$
By Conditions D.1, D.2, and D.3, this bound tends to zero. ∎
Lemma D.5.
Consider some general nuisance functions under $\mathbb{P}_{\infty}$, denoted by $\bar{S}_{\infty}$, $\bar{G}_{\infty}$, $\bar{\pi}_{\infty}$, and $q^{0}_{\infty}$. Then, $\mathbb{P}[\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\bar{S}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})}$
×$\quad\times\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})q^{0}_{\infty}(\mathbf{X})}{\bar{G}_{\infty}(y\mid a,\mathbf{X})\bar{\pi}_{\infty}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$
Proof.
We first express $\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})$ as
$\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\bar{S}_{\infty}(t\mid a,\mathbf{X})-\frac{q^{0}_{\infty}(\mathbf{X})}{\mathbb{P}(R=0)}\bar{S}_{\infty}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\bar{\pi}_{\infty}(a\mid\mathbf{X})}H_{\bar{S}_{\infty},\bar{G}_{\infty},t,a}(Y,\Delta,\mathbf{X}),$
where
$H_{\bar{S}_{\infty},\bar{G}_{\infty},t,a}(Y,\Delta,\mathbf{X})=\frac{\mathbb{I}(Y\leq t,\Delta=1)}{\bar{S}_{\infty}(Y\mid a,\mathbf{X})\bar{G}_{\infty}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\bar{\Lambda}_{\infty}(du\mid a,\mathbf{X})}{\bar{S}_{\infty}(u\mid a,\mathbf{X})\bar{G}_{\infty}(u\mid a,\mathbf{X})}.$
Following a result in Lemma 1 of Westling et al. [48], $\mathbb{E}\{H_{\bar{S}_{\infty},\bar{G}_{\infty},t,a}(Y,\Delta,\mathbf{X})\mid\mathbf{X}=\mathbf{x}\}$ equals
$-\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{x})\bar{G}(y\mid a,\mathbf{x})}{\bar{S}_{\infty}(y\mid a,\mathbf{x})\bar{G}_{\infty}(y\mid a,\mathbf{x})}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{x}).$
Therefore, it is straightforward that $\mathbb{P}[\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})]-\theta^{0}(t,a)$ equals
$\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\{\bar{S}_{\infty}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}\right]$
$+\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{q^{0}_{\infty}(\mathbf{X})}{q^{0}(\mathbf{X})}\bar{S}_{\infty}(t\mid a,\mathbf{X})\frac{\bar{\pi}(a\mid\mathbf{X})}{\bar{\pi}_{\infty}(a\mid\mathbf{X})}$
×$\quad\times\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})\bar{G}(y\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})\bar{G}_{\infty}(y\mid a,\mathbf{X})}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$
Furthermore, by Duhamel equation in Gill and Johansen [22], we have
$\bar{S}_{\infty}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})=-\bar{S}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y-\mid a,\mathbf{X})}(\Lambda^{0}_{\infty}-\Lambda^{0})(du\mid a,\mathbf{X}),$
for each $(t,a,\mathbf{x})$. Therefore, we further have $\mathbb{P}[\varphi^{\text{CCOD}}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\bar{S}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})}$
×$\quad\times\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})q^{0}_{\infty}(\mathbf{X})}{\bar{G}_{\infty}(y\mid a,\mathbf{X})\bar{\pi}_{\infty}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$ (5)
D.1.4 Proof of Theorem 2.8
Proof.
By the result in equation (D.1.3) with $\bar{\pi}_{\infty}=\bar{\pi}$, $\bar{G}_{\infty}=\bar{G}$, and $\bar{S}_{\infty}=\bar{S}$,
$\widehat{\theta}^{\text{CCOD}}_{n}(t,a)-\theta^{0}(t,a)$ $=\mathbb{P}_{n}[\varphi^{*\text{CCOD}}_{t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\varphi^{\text{CCOD}}_{t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\theta^{0}(t,a)\right].$
By Conditions D.1 and D.2, the second summand on the right-hand-side is $o_{p}(n^{-1/2})$ by Lemma D.4. By Lemma D.5, $\mathbb{P}[\widehat{\varphi}^{\text{CCOD}}_{t,a}]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\widehat{\bar{S}}_{m}(y\mid a,\mathbf{X})}$
×$\quad\times\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\widehat{\bar{\Lambda}}_{m}-\bar{\Lambda})(dy\mid a,\mathbf{X})\bigg{]}.$
By Duhamel equation in Gill and Johansen [22], we have that
$\frac{\bar{S}(y-\mid a,\mathbf{X})}{\widehat{\bar{S}}_{m}(y\mid a,\mathbf{X})}(\widehat{\bar{\Lambda}}_{m}-\bar{\Lambda})(dy\mid a,\mathbf{X})=\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X}),$
and so the above equals
$\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\right]$
$=\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\left\{\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}-1\right\}\int_{0}^{t}\frac{\bar{S}}{\widehat{\bar{S}}_{m}}(du\mid a,\mathbf{X})+\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})$
×$\qquad\times\int_{0}^{t}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})q^{0}(\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{]}$
$=\underbrace{\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})-\bar{\pi}(a\mid\mathbf{X})\}\{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\right]}_{I_{1}}$
$\quad+\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})$
×$\qquad\underbrace{\quad\times\int_{0}^{t}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\widehat{q}^{0}_{m}(\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})q^{0}(\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{]}}_{I_{2}}.$
We note that $I_{1}\leq\eta^{2}\mathbb{E}\left[|\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})-\bar{\pi}(a\mid\mathbf{X})|\cdot|\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})|\right]$. For $I_{2}$, we further expand it, similarly to the above process, as
$I_{2}$ $=\underbrace{\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\{\widehat{q}^{0}_{m}(\mathbf{X})-q^{0}(\mathbf{X})\}\{\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(t\mid a,\mathbf{X})\}\right]}_{I_{21}}$
$\quad+\mathbb{E}\bigg{[}\frac{\widehat{q}^{0}_{m}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\bar{\pi}(a\mid\mathbf{X})}{\widehat{\bar{\pi}}_{m}(a\mid\mathbf{X})}\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})$
×$\qquad\underbrace{\quad\times\int_{0}^{t}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\bigg{]}}_{I_{22}}.$
Then, we note that $I_{21}\leq\eta^{2}\mathbb{E}\left[|\widehat{q}^{0}_{m}(\mathbf{X})-q^{0}\mathbf{X})|\cdot|\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})-\bar{S}(y\mid a,\mathbf{X})|\right]$, and
$I_{22}\leq\eta^{2}\mathbb{E}\left[\widehat{\bar{S}}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})}{\widehat{\bar{G}}_{m}(y\mid a,\mathbf{X})}-1\right\}\left(\frac{\bar{S}}{\widehat{\bar{S}}_{m}}-1\right)(du\mid a,\mathbf{X})\right]$
by Condition D.2. Then, by notation in Condition D.3, $I_{1}+I_{2}\leq\eta^{2}\{\bar{r}_{n,t,a,1}+\bar{r}_{n,t,a,2}+\bar{r}_{n,t,a,3}\}$. Since $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}\leq 2$,
$\left|\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\theta^{0}(t,a)\right]\right|\leq 2\eta^{2}(1+\eta)\{\bar{r}_{n,t,a,1}+\bar{r}_{n,t,a,2}+\bar{r}_{n,t,a,3}\}=o_{p}(n^{-1/2}),$
by Condition D.3. This established the pointwise RAL property: $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{t,a})+o_{p}(n^{-1/2})$. Since $\varphi^{*\text{CCOD}}_{t,a}$ is uniformly bounded, $\mathbb{P}\{(\varphi^{*\text{CCOD}}_{t,a})^{2}\}<\infty$ and since $\mathbb{P}\{\varphi^{*\text{CCOD}}_{t,a}\}=0$, it follows that
$n^{1/2}\mathbb{P}_{n}(\widehat{\varphi}^{*\text{CCOD}}_{t,a})\to_{d}\mathcal{N}(0,\mathbb{P}\{(\varphi^{*\text{CCOD}}_{t,a})^{2}\}).$
For the uniform RAL, we use the same decomposition above. By conditions D.1, D.2 and D.3 and Lemma D.4,
$\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{\text{CCOD}}_{t,a}-\varphi^{\text{CCOD}}_{t,a}\right]\right|=o_{p}(n^{-1/2}).$
Therefore, we have that
$\sup_{u\in[0,t]}\left|\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}_{n}^{m}(\widehat{\varphi}^{\text{CCOD}}_{n,m,t,a}-\theta^{0}(t,a))\right|\leq\sup_{u\in[0,t]}2\eta^{2}\{\bar{r}_{n,u,a,1}+\bar{r}_{n,u,a,2}+\bar{r}_{n,u,a,3}\},$
which is $o_{p}(n^{-1/2})$ by Condition D.5. Thus, $\sup_{u\in[0,t]}\left|\widehat{\theta}^{\text{CCOD}}_{n}(u,a)-\theta^{0}(u,a)-\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{u,a})\right|=o_{p}(n^{-1/2})$. Since $\{\varphi^{*\text{CCOD}}_{u,a}:u\in[0,t]\}$ is a uniformly bounded $\mathbb{P}$-Donsker class by Lemma D.3, $\{n^{1/2}\mathbb{P}_{n}(\varphi^{*\text{CCOD}}_{u,a}):u\in[0,t]\}$ converges weakly to a tight mean-zero Gaussian process with covariance $(u,v)\mapsto\mathbb{P}(\varphi^{*\text{CCOD}}_{u,a}\varphi^{*\text{CCOD}}_{v,a})$. ∎
Remark D.6 (Double robustness of the CCOD estimator).
If we only focus on the consistency of $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$, then the condition $\bar{\pi}_{\infty}=\bar{\pi}$, $\bar{G}_{\infty}=\bar{G}$, $q^{0}_{\infty}=q^{0}$ and $\bar{S}_{\infty}=\bar{S}$ can be replaced by the following statement: For $\mathbb{P}$-almost all $\mathbf{x}$, there exist measurable sets $\bar{\mathcal{S}}_{x},\bar{\mathcal{G}}_{x}\subseteq[0,t]$ such that $\bar{\mathcal{S}}_{x}\cup\bar{\mathcal{G}}_{x}=[0,t]$ and $\bar{\Lambda}(u\mid a,\mathbf{x})=\bar{\Lambda}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\bar{\mathcal{S}}_{x}$ and $\bar{G}(u\mid a,\mathbf{x})=\bar{G}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\bar{\mathcal{G}}_{x}$. In addition, if $\bar{\mathcal{S}}_{x}$ is a strict subset of $[0,t]$, then $\bar{\pi}(a\mid\mathbf{x})=\bar{\pi}_{\infty}(a\mid\mathbf{x})$ and $q_{0}(\mathbf{x})=q^{0}_{\infty}(\mathbf{x})$ as well. Then, $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$ is consistent if Conditions D.1 and D.2 hold, and it is uniform consistent if Condition D.4 also holds.
This statement could be interpreted as that at a given time $t$, if either (i) the conditional survival model $\bar{S}$; or (ii) all other nuisance functions $\bar{G}$, $\bar{\pi}$ and $q^{0}$ are correctly specified (with other conditions above), $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$ is consistent.
The proof details of Remark D.6 are similar to those in the proof of Theorem 2 in Westling et al. [48], which are omitted here. A sketch is that by decomposing the integral $\int_{0}^{t}$ as $\int_{\bar{\mathcal{S}}_{x}}+\int_{\bar{\mathcal{S}}^{c}_{x}}$, where $\bar{\mathcal{S}}^{c}_{x}$ is the complement of set $\bar{\mathcal{S}}_{x}$, and $\bar{\mathcal{S}}_{x}^{c}\subseteq\bar{\mathcal{G}}_{x}$ by definition. Then it is straightforward to verify that when the statement in Remark D.6 holds, the following integral (also in the bias term (D.1.3))
$\int_{0}^{t}\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})q^{0}_{\infty}(\mathbf{X})}{\bar{G}_{\infty}(y\mid a,\mathbf{X})\bar{\pi}_{\infty}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})$
$=\left(\int_{\bar{\mathcal{S}}_{x}}+\int_{\bar{\mathcal{G}}_{x}}\right)\frac{\bar{S}(y-\mid a,\mathbf{X})}{\bar{S}_{\infty}(y\mid a,\mathbf{X})}\left\{\frac{\bar{G}(y\mid a,\mathbf{X})\bar{\pi}(a\mid\mathbf{X})q^{0}_{\infty}(\mathbf{X})}{\bar{G}_{\infty}(y\mid a,\mathbf{X})\bar{\pi}_{\infty}(a\mid\mathbf{X})q^{0}(\mathbf{X})}-1\right\}(\bar{\Lambda}_{\infty}-\bar{\Lambda})(dy\mid a,\mathbf{X})=0,$
which further implies $\mathbb{P}[\varphi^{\text{CCOD}}_{\infty,t,a}]-\theta^{0}(t,a)=0$. The remainder of proving the consistency or uniform consistency also needs Lemmata D.1D.3.
D.2 Theory of the source-site estimator
We first prove Theorem 2.10 about the EIF of the source-site estimator in the following Section D.2.1. We then provide proofs for the RAL and uniform RAL properties of the source-site estimator. The techniques employed parallel those used for the CCOD estimator, so some proofs are either omitted or provided with less detail.
D.2.1 Proof of Theorem 2.10
Proof.
Under Assumption 2.5, $\theta^{0}(t,a)=\mathbb{E}\{S^{0}(t\mid a,\mathbf{X})\mid R=0\}=\mathbb{E}\{S^{k}(t\mid a,\mathbf{X})\mid R=0\}$. Therefore,
00 0 $=\frac{\partial}{\partial\epsilon}\theta^{0}(t,a)\bigg{|}_{\epsilon=0}=\frac{\partial}{\partial\epsilon}\mathbb{E}\{S^{k}_{\epsilon}(t\mid a,\mathbf{X})\mid R=0\}\bigg{|}_{\epsilon=0}$
$=\mathbb{E}\{[S^{k}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}_{\mathbf{X}\mid R=0}\mid R=0\}+\mathbb{E}\left\{\int\frac{\partial}{\partial\epsilon}S^{k}_{\epsilon}(t\mid a,\mathbf{x})\bigg{|}_{\epsilon=0}\mu(d\mathbf{x})~{}\bigg{|}~{}R=0\right\}.$ (6)
Following a similar argument in the proof of Theorem 2.7 in Section D.1.1 (i.e., the way we derive the derivative for $\bar{S}_{\epsilon}$ wrt $\epsilon$ there), we can express the integrand of the second term above as
[…]
$=\iiint-\mathbb{I}(y\leq t,\delta=1)\frac{S^{k}(t\mid a,\mathbf{x})S^{k}(y-\mid a,\mathbf{x})}{S^{k}(y\mid a,\mathbf{x})D^{k}(y\mid\mathbf{x})}\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x},k)\mu(d\mathbf{x})$
$\quad+\iiiint\mathbb{I}(u\leq t,u\leq y)\frac{S^{k}(t\mid a,\mathbf{x})S^{k}(u-\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid\mathbf{x})}$
×˙$\quad\qquad\times\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x},k)N^{k}_{1}(du\mid a,\mathbf{x})\mu(d\mathbf{x})$
$=\iiint-\mathbb{I}(y\leq t,\delta=1)\frac{S^{k}(t\mid a,\mathbf{x})S^{k}(y-\mid a,\mathbf{x})}{S^{k}(y\mid a,\mathbf{x})D^{k}(y\mid\mathbf{x})}\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x},k)\mu(d\mathbf{x})$
$\quad+\iiint S^{k}(t\mid a,\mathbf{x})\int_{0}^{t\wedge y}\frac{S^{k}(u-\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid\mathbf{x})^{2}}N^{k}_{1}(du\mid a,\mathbf{x})$
×˙$\quad\qquad\times\dot{\ell}(y,\delta\mid a,\mathbf{x},k)\mathbb{P}(dy,d\delta\mid a,\mathbf{x},k)\mu(d\mathbf{x})$
$=\mathbb{E}\bigg{[}S^{k}(t\mid a,\mathbf{X})\frac{I(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left\{H^{k}(t\wedge Y,a,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid a,\mathbf{X})}{S^{k}(Y\mid a,\mathbf{X})D^{k}(Y\mid a,\mathbf{X})}\right\}$
×˙$\qquad\qquad\times\dot{\ell}(Y,\Delta\mid a,\mathbf{X},R=k)\bigg{]},$
where
$H^{k}(t,a,\mathbf{x})=\int_{0}^{t}\frac{S^{k}(u-\mid a,\mathbf{x})N_{1}^{k}(du\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid a,\mathbf{x})^{2}}.$
Now, we note that
$\mathbb{E}\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid A,\mathbf{X})}{S^{k}(Y\mid A,\mathbf{X})D^{k}(Y\mid A,\mathbf{X})}~{}\bigg{|}~{}A=a,\mathbf{X}=\mathbf{x},R=k\right]=\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{x})N_{1}^{k}(dy\mid a,\mathbf{x})}{S^{k}(y\mid a,\mathbf{x})D^{k}(y\mid a,\mathbf{x})},$
and
$\mathbb{E}\{H^{k}(t\wedge Y,A,\mathbf{X})\mid A=a,\mathbf{X}=\mathbf{x},R=k\}$
$=\iint^{t}\mathbb{I}(u\leq y)\frac{S^{k}(u-\mid a,\mathbf{x})N_{1}^{k}(du\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid a,\mathbf{x})^{2}}\mathbb{P}(dy\mid a,\mathbf{x},k)$
$=\int_{0}^{t}\mathbb{P}(Y\geq u\mid A=a,\mathbf{X}=\mathbf{x},R=k)\frac{S^{k}(u-\mid a,\mathbf{x})N_{1}^{k}(du\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid a,\mathbf{x})^{2}}\mathbb{P}(dy\mid a,\mathbf{x},k)$
$=\int_{0}^{t}\frac{S^{k}(u-\mid a,\mathbf{x})N_{1}^{k}(du\mid a,\mathbf{x})}{S^{k}(u\mid a,\mathbf{x})D^{k}(u\mid a,\mathbf{x})}.$
Therefore,
$\mathbb{E}\left[H^{k}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid A,\mathbf{X})}{S^{k}(Y\mid A,\mathbf{X})D^{k}(Y\mid A,\mathbf{X})}~{}\bigg{|}~{}A,\mathbf{X},R=k\right]=0$
almost surely. By properties of score functions and the tower property, the above implies that
[…]
$=\mathbb{E}\bigg{[}S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k\mid\mathbf{X})}\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}$
×$\qquad\times\left\{H^{k}(t\wedge Y,A,\mathbf{X})-\frac{\mathbb{I}(Y\leq t,\Delta=1)S^{k}(Y-\mid A,\mathbf{X})}{S^{k}(Y\mid A,\mathbf{X})D^{k}(Y\mid A,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})\bigg{]}.$
Combining these results with the facts that $N_{1}^{k}(du\mid a,\mathbf{x})/D^{k}(u\mid a,\mathbf{x})=\Lambda^{k}(du\mid a,\mathbf{x})$ and $D^{k}(u\mid a,\mathbf{x})=S^{k}(u-\mid x)G^{k}(u\mid a,\mathbf{x})$, we can rewrite (D.2.1) at the beginning as follows:
$\frac{\partial}{\partial\epsilon}\theta^{0}(t,a)\bigg{|}_{\epsilon=0}$
$=\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}[S^{k}(t\mid a,\mathbf{X})-\theta^{0}(t,a)]\dot{\ell}(\mathcal{O})-\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\mathbb{E}\bigg{\{}S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k\mid\mathbf{X})}$
×$\quad\quad\times\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left\{\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(y\mid\mathbf{X})G^{k}(y\mid a,\mathbf{X})}-\int_{0}^{t\wedge y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})~{}\bigg{|}~{}\mathbf{X}\bigg{\}}\bigg{]}$
$=\mathbb{E}\left[\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\{S^{k}(t\mid a,\mathbf{X})-\theta^{0}(t,a)\}\dot{\ell}(\mathcal{O})\right]-\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=0)}\frac{\mathbb{P}(R=0\mid\mathbf{X})}{\mathbb{P}(R=k\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})$
×$\quad\quad\times\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left\{\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(y\mid\mathbf{X})G^{k}(y\mid a,\mathbf{X})}-\int_{0}^{t\wedge y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right\}\dot{\ell}(\mathcal{O})\bigg{]}.$
Therefore, an uncentered EIF of $\theta^{0}(t,a)$ at $\mathbb{P}$ is found as
$\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})$
$=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}S^{0}(t\mid a,\mathbf{X})-\frac{\mathbb{I}(R=k)\mathbb{P}(R=0\mid\mathbf{X})}{\mathbb{P}(R=0)\mathbb{P}(R=k\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})$
×$\qquad\times\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(Y\mid a,\mathbf{X})G^{k}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid a,\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right].$
Observe that, by Bayes’s rule,
$\frac{\mathbb{P}(R=0\mid\mathbf{X})}{\mathbb{P}(R=k\mid\mathbf{X})}=\underbrace{\frac{\mathbb{P}(\mathbf{X}\mid R=0)}{\mathbb{P}(\mathbf{X}\mid R=k)}}_{\omega^{k,0}(\mathbf{X})}\cdot\frac{\mathbb{P}(R=0)}{\mathbb{P}(R=k)},$
where $\omega^{k,0}(\mathbf{X})$ is a density ratio term, we then have
$\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})$ $=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}S^{0}(t\mid a,\mathbf{X})-\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k)}\omega^{k,0}(\mathbf{X})S^{k}(t\mid a,\mathbf{X})$
×$\quad\quad\times\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(Y\mid a,\mathbf{X})G^{k}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid a,\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right].$
D.2.2 Regularity conditions and RAL of the source-site estimator
For site $R=k$, we denote $\pi^{k}$, $G^{k}$, $\omega^{k,0}$, $\Lambda^{k}$ and $S^{k}$ the truths of nuisance functions. We use $\pi_{\infty}^{k}$, $\omega^{k,0}_{\infty}$, $G_{\infty}^{k}$, $\Lambda_{\infty}^{k}$ and $S_{\infty}^{k}$ to denote some general probability limits for the nuisance function estimators.
Condition D.6.
There exist $\pi_{\infty}^{k}$, $\omega^{k,0}_{\infty}$, $G_{\infty}^{k}$, $\Lambda_{\infty}^{k}$ and $S_{\infty}^{k}$ such that
  1. (a)
    $\max_{m}\mathbb{P}\left[\frac{1}{\widehat{\pi}^{k}_{m}(a\mid\mathbf{X})}-\frac{1}{\pi_{\infty}^{k}(a\mid\mathbf{X})}\right]^{2}\to_{p}0$;
  2. (b)
    $\max_{m}\mathbb{P}\left[{\widehat{\omega}^{k,0}_{m}(\mathbf{X})}-{\omega^{k,0}_{\infty}(\mathbf{X})}\right]^{2}\to_{p}0$;
  3. (c)
    $\max_{m}\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{1}{\widehat{G}^{k}_{m}(u\mid a,\mathbf{X})}-\frac{1}{G_{\infty}^{k}(u\mid a,\mathbf{X})}\right|\right]^{2}\to_{p}0$;
  4. (d)
    $\max_{m}\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})}{\widehat{S}^{k}_{m}(u\mid a,\mathbf{X})}-\frac{S_{\infty}^{k}(t\mid a,\mathbf{X})}{S_{\infty}^{k}(u\mid a,\mathbf{X})}\right|\right]^{2}\to_{p}0$.
Condition D.7.
There exists an $\eta\in(0,\infty)$ such that for $\mathbb{P}$-almost all $\mathbf{x}$, $\widehat{\pi}^{k}_{m}(a\mid\mathbf{x})\geq 1/\eta$, $\pi^{k}_{\infty}(a\mid\mathbf{x})\geq 1/\eta$, $\widehat{\omega}^{k,0}_{m}(\mathbf{x})\leq\eta$, $\omega^{k,0}_{\infty}(\mathbf{x})\leq\eta$, $\widehat{G}^{k}_{m}(t\mid a,\mathbf{x})\geq 1/\eta$, and $G^{k}_{\infty}(t\mid a,\mathbf{x})\geq 1/\eta$ with probability tending to 1.
Condition D.8.
Define
$r_{n,t,a,1}^{k}$ $=\max_{m}\mathbb{P}\left|\{\widehat{\pi}^{k}_{m}(a\mid\mathbf{X})-\pi^{k}(a\mid\mathbf{X})\}\{\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})-S^{k}(t\mid a,\mathbf{X})\}\right|,$
$r_{n,t,a,2}^{k}$ $=\max_{m}\mathbb{P}\left|\{\widehat{\omega}^{k,0}_{m}(\mathbf{X})-\omega^{k,0}(\mathbf{X})\}\{\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})-S^{k}(t\mid a,\mathbf{X})\}\right|,\text{ and}$
$r_{n,t,a,3}^{k}$ $=\max_{m}\mathbb{P}\left|\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\left\{\frac{G^{k}(u\mid a,\mathbf{X})}{\widehat{G}^{k}_{m}(u\mid a,\mathbf{X})}-1\right\}\left(\frac{S^{k}}{\widehat{S}^{k}_{m}}-1\right)(du\mid a,\mathbf{X})\right|.$
Then, it holds that $r_{n,t,a,1}^{k}=o_{p}(n^{-1/2})$, $r_{n,t,a,2}^{k}=o_{p}(n^{-1/2})$ and $r_{n,t,a,3}^{k}=o_{p}(n^{-1/2})$.
Condition D.9.
$\max_{m}\mathbb{P}\left[\sup_{u\in[0,t]}\sup_{v\in[0,u]}\left|\frac{\widehat{S}^{k}_{m}(u\mid a,\mathbf{X})}{\widehat{S}^{k}_{m}(v\mid a,\mathbf{X})}-\frac{S_{\infty}^{k}(u\mid a,\mathbf{X})}{S_{\infty}^{k}(v\mid a,\mathbf{X})}\right|\right]^{2}\to_{p}0.$
Condition D.10.
It holds that $\sup_{u\in[0,t]}r_{n,u,a,1}^{k}=o_{p}(n^{-1/2})$, $\sup_{u\in[0,t]}r_{n,u,a,2}^{k}=o_{p}(n^{-1/2})$, and $\sup_{u\in[0,t]}r_{n,u,a,3}^{k}=o_{p}(n^{-1/2})$.
The following theorem formally states the (uniform) RAL of the source-site estimator.
Theorem D.7.
If Conditions D.6D.7 hold, with $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$, the CCOD holds, and Condition D.8 also holds, then $\widehat{\theta}^{k,0}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*k,0}_{t,a})+o_{p}(n^{-1/2})$. In particular, $n^{1/2}(\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a))$ then convergences in distribution to a normal random variable with mean zero and variance $\sigma^{2}=\mathbb{P}[(\varphi^{*k,0}_{t,a})^{2}]$. If Conditions D.9 and D.10 also hold, then
$\sup_{u\in[0,t]}\left|\widehat{\theta}^{k}_{n}(u,a)-\theta^{0}(u,a)-\mathbb{P}_{n}(\varphi^{*k,0}_{u,a})\right|=o_{p}(n^{-1/2}).$
In particular, $\{n^{1/2}(\widehat{\theta}^{k}_{n}(u,a)-\theta^{0}(u,a)):u\in[0,t]\}$ converges weakly as a process in the space $\ell^{\infty}([0,t])$ of uniformly bounded functions on $[0,t]$ to a tight mean zero Gaussian process with covariance function $(u,v)\mapsto\mathbb{P}(\varphi^{*k,0}_{u,a}\varphi^{*k,0}_{v,a})$.
To prove Theorem D.7, we first introduce and prove several lemmata in the next section.
D.2.3 Lemmata for the source-site estimator
To establish the RAL related results of $\widehat{\theta}_{n}^{k}(t,a)$, we start from considering the difference $\widehat{\theta}_{n}^{k}(t,a)-\theta^{0}(t,a)$. Recall that $\mathbb{P}_{n}^{m}$ is the empirical distribution corresponding to the $m$th validation set $\mathcal{V}_{m}$ from the entire data $\mathcal{O}$, and denote $\mathbb{G}_{n}^{m}$ the corresponding empirical process. Then, following the similar process used in (D.1.3), it can be shown that
$\widehat{\theta}_{n}^{k}(t,a)-\theta^{0}(t,a)$ $=\mathbb{P}_{n}[\varphi^{*k,0}_{\infty,t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{k,0}_{t,a}-\theta^{0}(t,a)\right].$ (7)
Next, similar to what we did for the CCOD estimator, in the following Lemma D.8, we establish the $L_{2}(\mathbb{P})$ norm distance (bound) between the estimated IF and the limiting IF for the source-site estimator.
Lemma D.8.
Under Condition D.7, there exists a universal constant $C=C(\eta)$ such that for each $k$, $m$, $n$, $t$, and $a$,
$\mathbb{P}[\widehat{\varphi}^{k,0}_{t,a}-\varphi^{k,0}_{\infty,t,a}]^{2}\leq C(\eta)\sum_{j=1}^{6}A^{k}_{j,n,m,t,a},$
where
$A^{k}_{1,n,m,t,a}$ $=\mathbb{P}\left[\frac{1}{\mathbb{P}_{n}^{m}(R=0)}-\frac{1}{\mathbb{P}(R=0)}\right]^{2},$
$A^{k}_{2,n,m,t,a}$ $=\mathbb{P}\left[\frac{1}{\mathbb{P}_{n}^{m}(R=k)}-\frac{1}{\mathbb{P}(R=k)}\right]^{2},$
$A^{k}_{3,n,m,t,a}$ $=\mathbb{P}\left[\widehat{\omega}^{k,0}_{m}(a\mid\mathbf{X})-\omega^{k,0}_{\infty}(a\mid\mathbf{X})\right]^{2},$
$A^{k}_{4,n,m,t,a}$ $=\mathbb{P}\left[\frac{1}{\widehat{\pi}^{k}_{m}(a\mid\mathbf{X})}-\frac{1}{\pi_{\infty}^{k}(a\mid\mathbf{X})}\right]^{2},$
$A^{k}_{5,n,m,t,a}$ $=\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{1}{\widehat{G}^{k}_{m}(u\mid a,\mathbf{X})}-\frac{1}{G_{\infty}^{k}(u\mid a,\mathbf{X})}\right|\right]^{2},$
$A^{k}_{6,n,m,t,a}$ $=\mathbb{P}\left[\sup_{u\in[0,t]}\left|\frac{\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})}{\widehat{S}^{k}_{m}(u\mid a,\mathbf{X})}-\frac{S_{\infty}^{k}(t\mid a,\mathbf{X})}{S_{\infty}^{k}(u\mid a,\mathbf{X})}\right|\right]^{2}.$
Proof.
We first denote
$B^{k}(\mathcal{V}_{m})$ $=\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}S^{k}(t\mid a,\mathbf{X})$
×$\quad\times\left[\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(Y\mid a,\mathbf{X})G^{k}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid a,\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\right],$
$C^{k}(\mathcal{V}_{m})$ $=B^{k}(\mathcal{V}_{m})\omega^{k,0}(\mathbf{X}).$
Then, we first have the following decomposition:
$\widehat{\varphi}^{k,0}_{t,a}-\varphi^{k,0}_{\infty,t,a}=\sum_{j=1}^{4}U^{k}_{j,n,m,t,a},$
where
$U^{k}_{1,n,m,t,a}$ $=\left[\frac{\mathbb{I}(R=0)}{\mathbb{P}_{n}^{m}(R=0)}-\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\right]\widehat{S}^{0}_{m}(t\mid a,\mathbf{x}),$
$U^{k}_{2,n,m,t,a}$ $=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\left[\widehat{S}^{0}_{m}(t\mid a,\mathbf{x})-S^{0}_{\infty}(t\mid a,\mathbf{x})\right],$
$U^{k}_{3,n,m,t,a}$ $=\left[\frac{\mathbb{I}(R=k)}{\mathbb{P}_{n}^{m}(R=k)}-\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k)}\right]\widehat{C}^{k}_{m}(\mathcal{V}_{m}),$
$U^{k}_{4,n,m,t,a}$ $=\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k)}\left[\widehat{C}^{k}_{m}(\mathcal{V}_{m})-C^{k}_{\infty}(\mathcal{V}_{m})\right].$
Now, for $U^{k}_{4,n,m,t,a}$, we further decompose it as
$U^{k}_{4,n,m,t,a}$ $=\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k)}\sum_{j=1}^{2}V^{k}_{j,n,m,t,a},$
where
$V^{k}_{1,n,m,t,a}$ $=B^{k}_{\infty}(\mathcal{V}_{m})\left[\widehat{\omega}^{k,0}_{m}(\mathbf{X})-\omega^{k,0}_{m}(\mathbf{X})\right],$
$V^{k}_{2,n,m,t,a}$ $=\widehat{\omega}^{k,0}_{m}(\mathbf{X})\left[\widehat{B}^{k}_{m}(\mathcal{V}_{m})-B^{k}_{\infty}(\mathcal{V}_{m})\right].$
The expression of $\widehat{B}^{k}_{m}(\mathcal{V}_{m})-B^{k}_{\infty}(\mathcal{V}_{m})$ is exactly the same as the Lemma 3 in Westling et al. [48], while we only need to replace the corresponding nuisance functions by the site-$k$ version here, so the detail is omitted. By the triangle inequality, we have $\mathbb{P}[\widehat{\varphi}^{k,0}_{t,a}-\varphi^{k,0}_{\infty,t,a}]^{2}\leq\left\{\sum_{j=1}^{4}\{\mathbb{P}[(U^{k}_{j,n,m,t,a})^{2}]\}^{1/2}\right\}^{2}$. Therefore, under Assumption 2.3 and Condition D.7, there exists a universal constant $C=C(\eta)$ such that the result in the statement holds. Thus, the proof is completed. ∎
Furthermore, we investigate how to make the empirical process term $\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a_{0}}\right]$ to be $o_{p}(n^{-1/2})$ to establish the asymptotic normality. We first recall Lemma D.2 in Section D.1.3 and related empirical process notation there. Then, we introduce the following Lemma D.9. For simplicity, $\omega$, $p^{0}$ and $p^{k}$ correspond to, respectively, $\omega^{k,0}$, $\mathbb{P}(R=0)$, $\mathbb{P}(R=k)$, and since $S^{0}=S^{k}$ is assumed, we just use $S$ to denote conditional survival of the event time for all $k$, and without loss of generality, $G=G^{k}$ and $\pi=\pi^{k}$ are also dependent to $R=k$.
Lemma D.9.
Let $S,\pi,\omega$, $p^{0}$, $p^{k}$ and $G$ be fixed, where $t\mapsto S(t\mid a,\mathbf{x})$ is assumed to be non-increasing for each $(a,\mathbf{x})$, and where $S(t_{0}\mid a,\mathbf{x})\geq 1/\eta$, $G(t_{0}\mid a,\mathbf{x})\geq 1/\eta$, $\pi(a_{0}\mid\mathbf{x})\geq 1/\eta$, $\omega(\mathbf{x})\leq\eta$, $p^{0}\geq 1/\eta$ and $p^{k}\geq 1/\eta$, for some $\eta\in(0,\infty)$. Then the class of influence functions $\mathcal{F}_{S,\pi,p^{0},q^{0},G,t_{0},a_{0}}=\{\varphi_{S,\pi,\omega,p^{0},p^{k},G,t_{0},a_{0}}:t\in[0,t_{0}]\}$ satisfies
$\sup_{\mathbb{Q}}N(\varepsilon\|F\|_{\mathbb{Q},2},\mathcal{F}_{S,\pi,\omega,p^{0},p^{k},G,t_{0},a_{0}},L_{2}(\mathbb{Q}))\leq 32/\varepsilon^{10},$
for any $\varepsilon\in(0,1]$ where $F=\eta(1+2\eta^{3})$ is an envelope of $\mathcal{F}_{S,\pi,\omega,p^{0},p^{k},G,t_{0},a_{0}}$, and the supremum is taken over all distributions $\mathbb{Q}$ on the sample space of the observed data.
Proof.
The class $\mathcal{F}_{S,\pi,\omega,p^{0},p^{k},G,t_{0},a_{0}}$ is uniformly bounded by $\eta(1+2\eta^{3})$ because of the assumed upper bounds of $1/p^{0},1/p^{k},\omega,1/\pi$ and $1/G$. Therefore, the envelop function can be taken as $F=\eta(1+2\eta^{3})$.
Define the following two functions $f_{t}$ and $h_{t}$ pointwise as
$f_{t}(\mathbf{x},r,a_{0},\delta,y)$ $=\frac{\mathbb{I}(r=k,a=a_{0},y\leq t,\delta=1)\omega(\mathbf{x})S(t\mid a_{0},\mathbf{x})}{p^{k}\pi(a_{0}\mid\mathbf{x})S(y\mid a_{0},\mathbf{x})G(y\mid a_{0},\mathbf{x})},$
$h_{t}(\mathbf{x},r,a_{0},y)$ $=\int\frac{\mathbb{I}(r=k,a=a_{0},u\leq y,u\leq t)\omega(\mathbf{x})S(t\mid a_{0},\mathbf{x})}{p^{k}\pi(a_{0}\mid\mathbf{x})S(u\mid a_{0},\mathbf{x})G(u\mid a_{0},\mathbf{x})}\Lambda(du\mid a_{0},\mathbf{x}).$
Consider classes $\mathcal{F}^{1}_{S,p^{0},t_{0},a_{0}}=\{(\mathbf{x},r)\mapsto\mathbb{I}(r=0)S(t\mid a_{0},\mathbf{x})/p^{0}:t\in[0,t_{0}]\}$, $\mathcal{F}^{2}_{S,p^{k},\omega,G,t_{0},a_{0}}=\{(\mathbf{x},r,a_{0},\delta,y)\mapsto f_{t}(\mathbf{x},r,a_{0},\delta,y):t\in[0,t_{0}]\}$, and $\mathcal{F}^{3}_{S,p^{k},\omega,G,t,a_{0}}=\{(\mathbf{x},r,a_{0},y)\mapsto h_{t}(\mathbf{x},r,a,y):t\in[0,t_{0}]\}$. We can then write
$\mathcal{F}_{S,\pi,\omega,p^{0},p^{k},G,t_{0},a_{0}}\subseteq\{f_{1}-f_{2}+f_{3}:f_{1}\in\mathcal{F}^{1}_{S,p^{0},t_{0},a_{0}},f_{2}\in\mathcal{F}^{2}_{S,p^{k},\omega,G,t_{0},a_{0}},f_{3}\in\mathcal{F}^{3}_{S,p^{k},\omega,G,t_{0},a_{0}}\}.$
Since $(\mathbf{x},r)\mapsto\mathbb{I}(r=0)S(t\mid a_{0},\mathbf{x})/p^{0}$ is non-increasing for all $(\mathbf{x},r)$ and uniformly bounded by $1/\eta$, Lemma D.2 implies that $\sup_{\mathbb{Q}}(\varepsilon,\mathcal{F}^{1}_{S,p^{0},t_{0},a_{0}},L_{2}(\mathbb{Q}))\leq 2\varepsilon^{2}$.
The $\mathcal{F}^{2}_{S,p^{k},\omega,G,t_{0},a_{0}}$ is contained in the product of the classes $\{y\mapsto\mathbb{I}(y\leq t):t\in[0,t_{0}]\}$, $\{\mathbf{x}\mapsto S(t\mid a_{0},\mathbf{x}):t\in[0,t_{0}]\}$, and the singleton class $\{(\mathbf{x},r,a_{0},\delta,y)\mapsto\mathbb{I}(r=k,a=a_{0},\delta=1,y\leq t)\omega(\mathbf{x})/[p^{k}\pi(a_{0}\mid\mathbf{x})S(y\mid a_{0},\mathbf{x})G(y\mid a_{0},\mathbf{x})]\}.$ The first two classes both have covering number $\sup_{\mathbb{Q}}N(\varepsilon,\cdot,L_{2}(\mathbb{Q}))\leq 2\varepsilon^{2}$ by Lemma D.2. The third class has uniform covering number 1 for all $\varepsilon$ because it can be covered with a single ball of any positive radius. In addition, $\mathcal{F}^{2}_{S,p^{k},\omega,G,t,a}$ is uniformly bounded by $\eta^{4}$. Therefore, $\sup_{\mathbb{Q}}N(\varepsilon\eta^{4},\mathcal{F}^{2}_{S,p^{k},\omega,G,t,a},L_{2}(\mathbb{Q}))\leq 4/\varepsilon^{4}$.
Furthermore, similar to the proof of Lemma D.3, we can rewrite $\mathcal{F}^{3}_{S,p^{k},\omega,G,t_{0},a_{0}}=\{(\mathbf{x},r,a_{0},y)\mapsto\int m_{t}(u,\mathbf{x},r,a_{0},y)\mu^{*}(du):t\in[0,t_{0}]\}$, where
$m_{t}(u,\mathbf{x},r,a_{0},y)=\frac{\mathbb{I}(r=k,a=a,u\leq y,u\leq t)\omega(\mathbf{x})S(t\mid a_{0},\mathbf{x})}{p^{k}\pi(a_{0}\mid\mathbf{x})S(u\mid a_{0},\mathbf{x})G(u\mid a_{0},\mathbf{x})}\lambda^{*}(u\mid a_{0},\mathbf{x}),$
with $\mu^{*}(du)$ is a dominating measure for $\mathbb{P}$-almost all $\mathbf{x}$ and $\lambda^{*}(du\mid a_{0},\mathbf{x})$ the Radon-Nikodym derivative of $\Lambda(\cdot\mid a_{0},\mathbf{x})$ with respect to $\mu^{*}$. Then, the class $\mathcal{M}_{t}=\{m_{t}:t\in[0,t_{0}]\}$ is contained in the product of the singleton class $\{(\mathbf{x},r,a_{0},y,u)\mapsto\mathbb{I}(r=k,a=a_{0},u\leq y)\omega(\mathbf{x})\lambda^{*}(u\mid a_{0},\mathbf{x})/[p^{k}\pi(a_{0}\mid\mathbf{x})S(u\mid a_{0},\mathbf{x})G(u\mid a_{0},\mathbf{x})]\}$ and the class $\{u\mapsto\mathbb{I}(u\leq t):t\in[0,t_{0}]\}$ and $\{\mathbf{x}\mapsto S(t\mid a_{0},\mathbf{x}):t\in[0,t]\}$, which as discussed above both have $L_{2}(\mathbb{Q})$ covering number bounded by $2/\varepsilon^{2}$ for any probability measure $\mathbb{Q}$. Therefore, $\sup_{\mathbb{Q}}N(\varepsilon\eta^{4},\mathcal{M}_{t},L_{2}(\mathbb{Q}))\leq 4/\varepsilon^{4}$. The remainder of the proof is the same as that of Lemma D.3 (except for replacing $\eta^{3}$ by $\eta^{4}$ here). ∎
Lemma D.10.
If Conditions D.6D.7 hold, $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}^{1/2}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a_{0}}\right]=o_{p}(n^{-1/2})$. If Conditions D.9 holds as well,
$\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{\infty,t,a_{0}}\right]\right|=o_{p}(n^{-1/2}).$
The proof of Lemma D.10 is similar to the proof of Lemma D.4, thus it is omitted.
Lemma D.11.
Consider some general nuisance functions under $\mathbb{P}_{\infty}$, denoted by $S^{0}_{\infty}$, $S^{k}_{\infty}$, $G^{k}_{\infty}$, $\pi^{k}_{\infty}$, and $\omega^{k,0}_{\infty}$ (equals 1 if $k=0$). Then, $\mathbb{P}[\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}S^{k}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})}$
×$\quad\times\left\{\frac{\omega^{k,0}_{\infty}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})\pi^{k}_{\infty}(a\mid\mathbf{X})}-1\right\}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}.$
Proof.
Following the first steps in the proof of Lemma D.5, we can show $\mathbb{P}[\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}\{S^{0}_{\infty}(t\mid a,\mathbf{X})-S^{0}(t\mid a,\mathbf{X})\}+\frac{q^{k}(\mathbf{X})}{\mathbb{P}(R=k)}\omega^{k,0}_{\infty}(\mathbf{X})S^{k}_{\infty}(t\mid a,\mathbf{X})\frac{\pi^{k}(a\mid\mathbf{X})}{\pi^{k}_{\infty}(a\mid\mathbf{X})}$
×$\qquad\qquad\times\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})G^{k}(y\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}$
$=\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\{S^{0}_{\infty}(t\mid a,\mathbf{X})-S^{0}(t\mid a,\mathbf{X})\}+\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\frac{\omega^{k,0}_{\infty}(\mathbf{X})}{\omega^{k,0}(\mathbf{X})}S^{k}_{\infty}(t\mid a,\mathbf{X})\frac{\pi^{k}(a\mid\mathbf{X})}{\pi^{k}_{\infty}(a\mid\mathbf{X})}$
×$\qquad\qquad\times\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})G^{k}(y\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}.$
In the second “$\mathbb{E}$” after “$=$”, we used the following relationship:
$\frac{q^{0}(\mathbf{X})}{q^{k}(\mathbf{X})}=\omega^{k,0}(\mathbf{X})\frac{\mathbb{P}(R=0)}{\mathbb{P}(R=k)}$
by Bayes’s rule. Furthermore, by Duhamel equation in Gill and Johansen [22] used in Lemma D.5 again, $\mathbb{P}[\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}S^{k}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})}$
×$\qquad\times\left\{\frac{\omega^{k,0}_{\infty}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})\pi^{k}_{\infty}(a\mid\mathbf{X})}-1\right\}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}.$ (8)
D.2.4 Proof of Theorem D.7
Proof.
By (D.2.3) with $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$,
$\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a)$ $=\mathbb{P}_{n}[\varphi^{*k,0}_{t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{k,0}_{t,a}-\theta^{0}(t,a)\right].$
By Conditions D.6 and D.7, the second summand on the right-hand-side is $o_{p}(n^{-1/2})$ by Lemma D.10. By Lemma D.11, $\mathbb{P}[\widehat{\varphi}^{k,0}_{t,a}]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{\widehat{S}^{k}_{m}(y\mid a,\mathbf{X})}$
×$\qquad\times\left\{\frac{\widehat{\omega}^{k,0}_{m}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})\widehat{G}^{k}_{m}(y\mid a,\mathbf{X})\widehat{\pi}^{k}_{m}(a\mid\mathbf{X})}-1\right\}(\widehat{\Lambda}^{k}_{m}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}.$
By a similar decomposition in the proof of Theorem 2.8 in Section D.1.4 using Duhamel equation in Gill and Johansen [22], and by notation in Condition D.8, we find that the above bias term can be bounded by $\eta^{2}\{r_{n,t,a,1}^{k}+r_{n,t,a,2}^{k}+r_{n,t,a,3}^{k}\}$ over $m$. Since $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}\leq 2$, we have
$\left|\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{k,0}_{t,a}-\theta^{0}(t,a)\right]\right|\leq 2\eta^{2}\left\{r_{n,t,a,1}^{k}+r_{n,t,a,2}^{k}+r_{n,t,a,3}^{k}\right\}=o_{p}(n^{-1/2}),$
by Condition D.8. This established the pointwise RAL property: $\widehat{\theta}^{k,0}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*k,0}_{t,a})+o_{p}(n^{-1/2})$. Since $\varphi^{*k,0}_{t,a}$ is uniformly bounded, $\mathbb{P}\{(\varphi^{*k,0}_{t,a})^{2}\}<\infty$ and since $\mathbb{P}\{\varphi^{*k,0}_{t,a}\}=0$, it follows that
$n^{1/2}\mathbb{P}_{n}(\widehat{\varphi}^{*k,0}_{t,a})\to_{d}\mathcal{N}(0,\mathbb{P}\{(\varphi^{*k,0}_{t,a})^{2}\}).$
For the uniform RAL, by conditions D.6, D.7 and D.8 and Lemma D.10,
$\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\sup_{u\in[0,t]}\left|\mathbb{G}_{n}^{m}(\widehat{\varphi}^{k,0}_{t,a}-\varphi^{k,0}_{t,a})\right|=o_{p}(n^{-1/2}).$
Therefore, we have that
$\sup_{u\in[0,t]}\left|\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}_{n}^{m}(\widehat{\varphi}^{k,0}_{n,m,t,a}-\theta^{0}(t,a))\right|\leq\sup_{u\in[0,t]}2\eta^{2}\{r_{n,t,a,1}^{k}+r_{n,t,a,2}^{k}+r_{n,t,a,3}^{k}\},$
which is $o_{p}(n^{-1/2})$ by Condition D.10. Therefore, $\sup_{u\in[0,t]}\left|\widehat{\theta}^{k}_{n}(u,a)-\theta^{0}(u,a)-\mathbb{P}_{n}(\varphi^{*k,0}_{u,a})\right|=o_{p}(n^{-1/2})$. Since $\{\varphi^{*k,0}_{u,a}:u\in[0,t]\}$ is a uniformly bounded $\mathbb{P}$-Donsker class by Lemma D.9, $\{n^{1/2}\mathbb{P}_{n}\{\varphi^{*k,0}_{u,a}\}:u\in[0,t]\}$ converges weakly to a tight mean-zero Gaussian process with covariance $(u,v)\mapsto\mathbb{P}(\varphi^{*k,0}_{u,a}\varphi^{*k,0}_{v,a})$. ∎
Remark D.12 (Double robustness of the source-site estimator).
If we only need the consistency of $\widehat{\theta}_{n}^{k}(t,a)$, then condition $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$ can be replaced by the following statement: For $\mathbb{P}$-almost all $\mathbf{x}$, there exist measurable sets $\mathcal{S}^{k}_{x},\mathcal{G}^{k}_{x}\subseteq[0,t]$ such that $\mathcal{S}^{k}_{x}\cup\mathcal{G}^{k}_{x}=[0,t]$ and $\Lambda^{k}(u\mid a,\mathbf{x})=\Lambda_{\infty}^{k}(u\mid a,\mathbf{x})$ for all $u\in S^{k}_{x}$ and $G(u\mid a,\mathbf{x})=G^{k}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\mathcal{G}^{k}_{x}$. In addition, if $\mathcal{S}^{k}_{x}$ is a strict subset of $[0,t]$, then $\pi^{k}(a\mid\mathbf{x})=\pi^{k}_{\infty}(a\mid\mathbf{x})$ and $\omega^{k,0}(\mathbf{x})=\omega^{k,0}_{\infty}(\mathbf{x})$ as well. Then, $\widehat{\theta}_{n}^{k}(t,a)$ is consistent if Conditions D.6 and D.7 hold, and it is uniform consistent if Condition D.9 also holds. This statement could be interpreted as that at a given time $t$, if either (i) the conditional survival model $S^{k}$; or (ii) all other nuisance functions $G^{k}$, $\pi^{k}$ and $\omega^{k,0}$ are correctly specified (with other conditions above), $\widehat{\theta}_{n}^{k}(t,a)$ is consistent.
The proof sketch of Remark D.12 is similar to that under Remark D.6 via examining the bias term (D.2.3), which is omitted here.
D.3 Theory for the federated estimator
In this section, we present the theoretical properties of the federated estimator. Given that our proposed weights, $\boldsymbol{\eta}_{t,a}$, are both time- and treatment-specific, we focus on the pointwise convergence properties of the federated estimator.
Let the set of all source site indices be $\mathcal{S}={1,\dots,K-1}$. We then define the oracle selection space for $\boldsymbol{\eta}_{t,a}$, and the corresponding weight space as:
$\mathcal{S}^{*}_{t,a}=\{k\in\mathcal{S}:\theta^{k}(t,a)=\theta^{0}(t,a)\},~{}~{}\text{and}~{}~{}\mathbb{R}^{S^{*}_{t,a}}=\{\boldsymbol{\eta}_{t,a}\in\mathbb{R}^{K-1}:\eta^{j}_{t,a}=0,\forall j\not\in\mathcal{S}^{*}_{t,a}\},$
respectively.
The space $\mathcal{S}^{*}_{t,a}$ is both time- and treatment-varying, indicating that a source site may not consistently be useful or unhelpful across different time points or treatments. However, it offers the advantage of increased flexibility and adaptivity, allowing for more effective borrowing of information at different points along the survival functions. Based on the theory presented in Section D.2, for $k\in\mathcal{S}^{*}_{t,a}$, the site-specific estimator $\widehat{\theta}^{k,0}_{n}(t,a)$ is consistent for $\theta^{0}(t,a)$ for any given $t\in[0,\tau]$ and $a\in\{0,1\}$.
We begin by assuming fixed $\boldsymbol{\eta}_{t,a}=(\eta_{t,a}^{0},\eta_{t,a}^{1},\dots,\eta_{t,a}^{K-1})$. We invoke Lemmata 4 and 5 in Han et al. [25], which state that the proposed adaptive estimation for $\eta_{t,a}^{k}$ as shown in (2) allows for (i) the recovery of the optimal $\bar{\eta}^{k}_{t,a}$ by the estimator $\widehat{\eta}^{k}_{t,a}$, and (ii) the uncertainty induced by $\widehat{\eta}^{k}_{t,a}$ is negligible when estimating $\theta^{0}(t,a)$. We require regularity Conditions D.6, D.7 and D.8 for the pointwise convergence result in Theorem D.7 hold. Let us denote the federated estimator by plugging-in the fixed $\boldsymbol{\eta}_{t,a}$ as
$\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})=\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\widehat{\theta}^{0}_{n}(t,a)+\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\widehat{\theta}^{k,0}_{n}(t,a).$
Recall that notation $\mathcal{H}_{t,a}$ defined in (1):
$\mathcal{H}_{t,a}(\mathcal{O};S,G)=\frac{\mathbb{I}(Y\leq t,\delta=1)}{S(Y\mid a,\mathbf{X})G(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda(du\mid a,\mathbf{X})}{S(u\mid a,\mathbf{X})G(u\mid a,\mathbf{X})}.$
Let us then write
$\xi^{0,(1)}(\mathcal{O})$ $=S^{0}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{0}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};S^{0},\Lambda^{0},G^{0}),$
$\xi^{k,0,(1)}(\mathcal{O})$ $=\omega^{k,0}(\mathbf{X})S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};S^{k},\Lambda^{k},G^{k}),$
$\xi^{0,(2)}(\mathcal{O})$ $=S^{0}(t\mid a,\mathbf{X})-\theta^{0}(t,a),$
and $n_{k}=\sum_{i=1}^{n}\mathbb{I}(R_{i}=k)$ for $k=0,1,\dots,K-1$.
Then,
$\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})-\theta^{0}(t,a)$
$=\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\left\{\widehat{\theta}^{0}_{n}(t,a)-\theta^{0}(t,a)\right\}+\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\left\{\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a)\right\}$
$=\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\frac{1}{n_{0}}\sum_{i=1}^{n}\mathbb{I}(R_{i}=0)\left\{\widehat{\xi}^{0,(2)}(\mathcal{O}_{i})-\widehat{\xi}^{0,(1)}(\mathcal{O}_{i})\right\}$
$\quad+\sum_{k\in\mathcal{S}}\frac{1}{n_{0}}\sum_{i=1}^{n}\mathbb{I}(R_{i}=0)\eta^{k}_{t,a}\widehat{\xi}^{0,(2)}(\mathcal{O}_{i})-\sum_{k\in\mathcal{S}}\frac{1}{n_{k}}\sum_{i=1}^{n}\mathbb{I}(R_{i}=k)\eta^{k}_{t,a}\widehat{\xi}^{k,0,(1)}(\mathcal{O}_{i})$
$=\frac{1}{n}\sum_{i=1}^{n}\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\mathbb{I}(R_{i}=0)\frac{\widehat{\xi}^{0,(2)}(\mathcal{O}_{i})-\widehat{\xi}^{0,(1)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=0)}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(R_{i}=0)\left(\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\frac{\widehat{\xi}^{0,(2)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=0)}-\frac{1}{n}\sum_{k\in\mathcal{S}}\sum_{i=1}^{n}\mathbb{I}(R_{i}=k)\eta^{k}_{t,a}\frac{\widehat{\xi}^{k,0,(1)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=k)}.$ (9)
The asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})$ equals the variance of the influence function of (D.3). Let us denote it as $\mathcal{V}_{t,a}=\mathcal{V}_{t,a}(\boldsymbol{\eta}_{t,a})$. We highlight its dependence to the federated weights vector $\boldsymbol{\eta}_{t,a}$ here because in the below (11), we consider an optimization program for deriving the weights based on minimizing the (estimated) asymptotic variance.
Under the assumption of i.i.d. participants within each site, we have
$\mathcal{V}_{t,a}$ $=\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)^{2}\frac{\mathbb{V}\{\xi^{0,(2)}(\mathcal{O}_{i})-\xi^{0,(1)}(\mathcal{O}_{i})\mid R_{i}=0\}}{\mathbb{P}(R_{i}=0)}$
$\quad+\left(\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)^{2}\frac{\mathbb{V}\{\xi^{0,(2)}(\mathcal{O}_{i})\mid R_{i}=0\}}{\mathbb{P}(R_{i}=0)}$
$\quad+2\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\left(\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\frac{\text{Cov}\{\xi^{0,(2)}(\mathcal{O}_{i})-\xi^{0,(1)}(\mathcal{O}_{i}),\xi^{0,(2)}(\mathcal{O}_{i})\mid R_{i}=0\}}{\mathbb{P}(R_{i}=0)}$
$\quad+\sum_{k\in\mathcal{S}}(\eta^{k}_{t,a})^{2}\frac{\mathbb{V}\{\xi^{k,0,(1)}(\mathcal{O}_{i})\mid R_{i}=k\}}{\mathbb{P}(R_{i}=k)}.$ (10)
With appropriate boundedness conditions on variance and covariance of the influence functions, this variance is finite. Consequently, the asymptotic distribution of $\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})$ is expressed as
$\sqrt{n}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,\mathcal{V}_{t,a}).$
We further define the optimal adaptive weights $\bar{\boldsymbol{\eta}}_{t,a}$ as follows:
$\bar{\boldsymbol{\eta}}_{t,a}=\underset{\eta^{k}_{t,a}=0,\forall k\not\in\mathcal{S}^{*}_{t,a}}{\text{arg min}}\mathcal{V}_{t,a}(\boldsymbol{\eta}_{t,a}).$ (11)
By leveraging Lemmata 4 and 5 from Han et al. [25], we can recover the optimal weights $\bar{\boldsymbol{\eta}}_{t,a}$ with negligible uncertainty for estimating $\theta^{0}(t,a)$ if we estimate ${\boldsymbol{\eta}}_{t,a}$ using (2), akin to adaptive Lasso [57, 18]. The consistency of $\widehat{\mathcal{V}}_{t,a}=\widehat{\mathcal{V}}_{t,a}(\widehat{\boldsymbol{\eta}}_{t,a})$ follows when we can effectively approximate ${\mathcal{V}}_{t,a}(\bar{\boldsymbol{\eta}}_{t,a})$ with $\widehat{\mathcal{V}}_{t,a}$. Thus,
$\sqrt{n/\widehat{\mathcal{V}}_{t,a}}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a)-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,1).$
We now proceed to analyze the efficiency gain resulting from the federation process. The estimator relies only on the target data is denoted as $\widehat{\theta}^{0}_{n}(t,a)=\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}^{0}_{t,a})$, where $\boldsymbol{\eta}^{0}_{t,a}$ assigns all weights to the target and none to the source. In contrast, the estimator that leverages the proposed adaptive ensemble approach is denoted as $\widehat{\theta}^{\text{fed}}_{n}(t,a;\widehat{\boldsymbol{\eta}}_{t,a})$. Here $\widehat{\boldsymbol{\eta}}_{t,a}$ can recover the optimal weights $\bar{\boldsymbol{\eta}}_{t,a}$ that are associated with the minimum asymptotic variance. Consequently, the variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a;\widehat{\boldsymbol{\eta}}_{t,a})$ is no larger than that of the estimator relying solely on the target data since $\boldsymbol{\eta}^{0}_{t,a}$ is generally not the variance minimizer.
To establish that the asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a;\widehat{\boldsymbol{\eta}}_{t,a})$ is strictly smaller than that of the estimator based solely on the target data $\widehat{\theta}^{0}_{n}(t,a)$, we adopt Proposition 1 in Han et al. [25] with a modified informative source condition (modified Assumption 3(b) in Han et al. [25]).
Specifically, for each source site $s\in\mathcal{S}^{*}_{t,a}$, we define $\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})$ a federated estimator where $\eta^{s}_{t,a}$ is the optimal ensemble weight of site $s$ if we only consider target site and this source site $s$ for the federation. Then, the modified informative source condition is given as
$\left|\text{Cov}\left[\sqrt{n}\widehat{\theta}^{0}_{n}(t,a),\sqrt{n}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)\right\}\right]\right|\geq\varepsilon,$
for some $\varepsilon>0$, where $\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)$ can be expressed as
$\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)$
=$=\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\theta^{0}(t,a)\right\}-\left\{\widehat{\theta}^{0}_{n}(t,a)-\theta^{0}(t,a)\right\}$
$=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(R_{i}=0)(1-\eta^{s}_{t,a})\frac{\widehat{\xi}^{0,(2)}(\mathcal{O}_{i})-\widehat{\xi}^{0,(1)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=0)}+\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(R_{i}=0)\eta^{s}_{t,a}\frac{\widehat{\xi}^{0,(2)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=0)}$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(R_{i}=s)\eta^{s}_{t,a}\frac{\widehat{\xi}^{s,0,(1)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=s)}-\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(R_{i}=0)\frac{\widehat{\xi}^{0,(2)}(\mathcal{O}_{i})-\widehat{\xi}^{0,(1)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=0)}$
$=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(R_{i}=0)\eta^{s}_{t,a}\frac{\widehat{\xi}^{0,(1)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=0)}-\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(R_{i}=s)\eta^{s}_{t,a}\frac{\widehat{\xi}^{s,0,(1)}(\mathcal{O}_{i})}{\widehat{\mathbb{P}}(R_{i}=s)}.$
Therefore, it is straightforward to see that the modified condition can be achieved if $\eta^{s}_{t,a}>0$.